### v1

# The FFT Strikes Back: An Efficient Alternative to Self-Attention

Abstract
Conventional self-attention mechanisms incur quadratic complexity, limiting their scalability on long sequences. We introduce FFTNet, an adaptive spectral filtering framework that leverages the Fast Fourier Transform (FFT) to achieve global token mixing in $\mathcal{O}(n\log n)$ time. By transforming inputs into the frequency domain, FFTNet exploits the orthogonality and energy preservation guaranteed by Parseval’s theorem to capture long-range dependencies efficiently. A learnable spectral filter and modReLU activation dynamically emphasize salient frequency components, providing a rigorous and adaptive alternative to traditional self-attention. Experiments on the Long Range Arena and ImageNet benchmarks validate our theoretical insights and demonstrate superior performance over fixed Fourier and standard attention models.
1 Introduction
Conventional self-attention mechanisms capture global interactions through explicit pairwise computations, which results in a quadratic computational complexity that can be prohibitive for long sequences. In contrast, our work introduces an adaptive spectral filtering framework that leverages the Fast Fourier Transform (FFT) to perform global token mixing with a mathematically elegant and scalable approach.
Our method begins by transforming the input sequence into the frequency domain, where orthogonal frequency components naturally encode long-range dependencies. This not only reduces the computational complexity to $\mathcal{O}(n\log n)$ but also preserves the energy of the original signal, as ensured by Parseval’s theorem. Such a transformation facilitates efficient global interactions without the need for exhaustive pairwise comparisons.
A central innovation of our framework is the integration of a learnable spectral filter. This adaptive component modulates the Fourier coefficients based on a global context vector, enabling the model to dynamically emphasize salient frequency bands that are critical for capturing complex patterns. Furthermore, the application of nonlinear activations to both the real and imaginary parts of the filtered signal enhances the model’s expressivity, allowing it to represent higher-order interactions that go beyond the scope of linear transformations.
In essence, our adaptive spectral filtering framework combines the computational efficiency of FFT-based transformations with adaptive, context-sensitive filtering and nonlinear processing. This synthesis offers a rigorous and expressive alternative to traditional self-attention, providing a robust solution for modeling long-range dependencies in sequence data.
2 Related Work
In this section, we review existing methods aimed at improving the efficiency of sequence models. We first discuss the complexity issues inherent in self-attention (Section 2.1), then highlight Fourier-based approaches (Section 2.2) and other approximation techniques (Section 2.3). We then examine orthogonal matrix decomposition methods (Section 2.4), and finally position our adaptive spectral filtering method within this landscape (Section 2.5).
2.1 Self-Attention Complexity
The original Transformer architecture Vaswani et al. (2017) uses pairwise dot-product attention, incurring a computational and memory cost of $\mathcal{O}(n^{2})$, where $n$ is the sequence length. As $n$ grows, this quadratic complexity quickly becomes infeasible for long sequences in tasks such as language modeling and long-context document understanding.
2.2 Fourier-Based Mixing
Fourier-based approaches leverage the Fast Fourier Transform (FFT) Cooley and Tukey (1965) to achieve more efficient global mixing of tokens. FNet Lee-Thorp et al. (2022), for example, replaces the self-attention sublayer with a fixed Fourier transform, drastically lowering computational overhead. However, the use of a static transform limits its capacity to adapt to varying inputs or highlight task-specific frequency components.
2.3 Linear, Sparse, and Low-Rank Approximations
Beyond Fourier methods, several alternative strategies aim to reduce the cost of self-attention. Performer Choromanski et al. (2021) and linear transformer variants Katharopoulos et al. (2020) approximate the softmax attention matrix to achieve linear or near-linear complexity. Meanwhile, Reformer Kitaev et al. (2020), Linformer Wang et al. (2020), and BigBird Zaheer et al. (2020) employ sparse or low-rank approximations, extending the effective context length without paying the full quadratic price. Other approaches like Synthesizer Tay et al. (2020) and MLP-Mixer Tolstikhin et al. (2021) avoid explicit token-pair interactions, replacing them with fixed or learned mixing operations.
2.4 Orthogonal Matrix Decomposition Methods
Orthogonal (or unitary) transformations provide a powerful avenue for stable and efficient sequence modeling. A key advantage of orthogonal decompositions is their norm-preserving property, which can mitigate issues such as vanishing or exploding gradients Wisdom et al. (2016). In the context of RNNs, unitary or orthonormal recurrent weights have been shown to preserve long-term dependencies while keeping representations stable Arjovsky et al. (2016); Lezcano-Casado and Martínez-Rubio (2019). From another perspective, the discrete Fourier transform (DFT) itself is an orthonormal transformation (up to scaling) that can mix tokens globally without explicit pairwise attention.
The FFT-based approaches discussed above can be viewed as a special class of such orthonormal transforms, where the matrix is structured by the DFT. More general orthogonal transformations—whether learned or hand-crafted—have also been proposed to reduce complexity or enhance stability in modern architectures. These include fast variants of orthonormal transforms, often parameterized in ways that ensure orthogonality is preserved throughout training Lezcano-Casado and Martínez-Rubio (2019). Within Transformers, adopting orthogonal or unitary blocks has been explored to stabilize training and capture global structure, although these methods may not always achieve the same $\mathcal{O}(n\log n)$ cost as the FFT. Nonetheless, they highlight a broad paradigm wherein structured or parameterized orthonormal decompositions serve as efficient global mixing mechanisms.
2.5 Adaptive Spectral Filtering in Context
Our work diverges from both fixed Fourier-based schemes and the various attention approximations by incorporating a learnable filter in the frequency domain. This adaptive mechanism leverages the theoretical underpinnings of FFT-based transformations—including energy preservation via Parseval’s theorem—while permitting dynamic reweighting of salient frequency bands. Thus, our method maintains an $\mathcal{O}(n\log n)$ complexity yet provides richer expressivity than fixed spectral mixing approaches. In contrast to purely approximate or sparse attention mechanisms, adaptive spectral filtering offers a direct and theoretically grounded route to capture long-range dependencies efficiently.
Overall, while approaches such as FNet, Performer, and sparse transformers demonstrate that either fixed or approximate token mixing can reduce computational overhead, our adaptive spectral filtering strategy uniquely merges the efficiency of the FFT with a learnable, input-dependent spectral filter. This provides a compelling combination of scalability and adaptability, which is crucial for complex sequence modeling tasks.
3 Adaptive Spectral Filtering Method
In this section, we present an adaptive spectral filtering framework that eliminates the need for explicit pairwise interactions in global token mixing. Instead of relying on dot-product-based self-attention, we employ the discrete Fourier transform (DFT) to capture long-range dependencies efficiently. By adaptively modulating the resulting frequency components and then applying an inverse transform, our method strikes an effective balance between expressive power and computational cost, supported by strong theoretical guarantees.
3.1 Motivation
Standard attention mechanisms compute pairwise interactions between tokens, incurring a quadratic cost in sequence length. As the number of tokens grows, this approach rapidly becomes prohibitive. In contrast, the Fourier transform decomposes a sequence into frequency components, inherently encoding global interactions in $\mathcal{O}(n\log n)$ time. To enhance representational power, we introduce a learnable filter in the frequency domain, enabling the model to emphasize salient frequency bands while retaining computational efficiency.
3.2 Method Description
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be the input sequence of length $n$ and embedding dimension $d$. Our method comprises four steps:
1. Fourier Transform.
We first apply the discrete Fourier transform along the token dimension:
$\mathbf{F}=\operatorname{FFT}(\mathbf{X})\;\in\;\mathbb{C}^{n\times d}.$
This operation represents each embedding across orthogonal frequency components, enabling global interactions without explicit pairwise comparisons.
2. Adaptive Spectral Filtering.
To selectively emphasize important frequencies, we use a learnable filter. First, compute a global context vector:
$\mathbf{c}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{X}_{i},$
and pass it through a multi-layer perceptron (MLP) to obtain a modulation tensor:
$\Delta\mathbf{W}=\operatorname{MLP}(\mathbf{c})\;\in\;\mathbb{R}^{n\times d}.$
We define the final filter as
$\mathbf{W}=\mathbf{W}_{\mathrm{base}}+\Delta\mathbf{W},$
where $\mathbf{W}_{\mathrm{base}}$ is a fixed base filter (often initialized to all ones). The adaptive filtering step is then
$\tilde{\mathbf{F}}=\mathbf{F}\odot\mathbf{W},$
which reweights the Fourier coefficients element-wise according to the global context.
3. Nonlinear Activation (modReLU).
To capture higher-order relationships in the complex frequency domain, we apply the modReLU Arjovsky et al. (2016) activation, defined for a complex number $z=re^{i\theta}$ (with $r=|z|$ and $\theta=\arg(z)$) as:
$\operatorname{modReLU}(z)\;=\;\begin{cases}(r+b)\,e^{i\theta},&\text{if }r+b>0,\\[6.0pt] 0,&\text{otherwise},\end{cases}$
where $b$ is a learnable bias. This operation applies a ReLU-like threshold to the magnitude while preserving the phase, making it well suited for frequency-domain representations. The element-wise modReLU is then:
$\tilde{\mathbf{F}}=\operatorname{modReLU}\bigl{(}\tilde{\mathbf{F}}\bigr{)}.$
4. Inverse Fourier Transform.
Finally, we return to the token domain via the inverse Fourier transform, retaining the real component of the reconstructed signal:
$\mathbf{Y}=\operatorname{IFFT}\bigl{(}\tilde{\mathbf{F}}\bigr{)}\;\in\;\mathbb{R}^{n\times d}.$
This yields a globally mixed representation that incorporates adaptive filtering and nonlinear transformations in the frequency domain.
3.3 Theoretical Justification for FFT over Self-Attention
Our FFT-based adaptive spectral filtering approach offers several advantages over conventional self-attention:
  • Efficient Global Mixing: By decomposing inputs into frequency components, the FFT provides global interactions in $\mathcal{O}(n\log n)$ time—much more scalable than the $\mathcal{O}(n^{2})$ of self-attention.
  • Implicit, Adaptive Attention: The adaptive spectral filter effectively learns a frequency-domain mask informed by the global context vector. This mask reweights crucial frequency bands similarly to attention weights but avoids explicit pairwise computations.
  • Greater Expressivity via Nonlinearity: While the Fourier transform itself is linear, applying modReLU to the complex coefficients enriches representational capacity. This allows the model to capture intricate, higher-order patterns that purely linear operations may miss.
  • Energy Preservation and Stability: Parseval’s theorem ensures the input signal’s norm (energy) is preserved by the FFT. Thus, our method maintains stability by avoiding accidental loss of crucial information.
By leveraging frequency-domain global mixing, learnable spectral filtering, and the modReLU nonlinear activation, our method is both theoretically grounded and computationally efficient, serving as a robust alternative to self-attention.
3.4 Computational Complexity
The central operations in our method are the Fast Fourier Transform (FFT) and its inverse, each costing $\mathcal{O}(n\log n)$ per channel. In contrast, self-attention requires $\mathcal{O}(n^{2})$ time for pairwise computations. The adaptive filtering and activation steps introduce only a linear overhead $\mathcal{O}(n)$, keeping the overall complexity at $\mathcal{O}(n\log n)$.
3.5 Summary
In essence, adaptive spectral filtering transforms the input into the frequency domain, applies a learnable nonlinear modulation (modReLU), and then inverts the transform to obtain a globally mixed representation. This provides:
  • A favorable $\mathcal{O}(n\log n)$ complexity.
  • Robust mechanisms for modeling long-range dependencies.
  • An efficient, expressive alternative to self-attention.
The code is available here: https://github.com/jacobfa/fft.
3.6 Proof of Computational Complexity
Fourier and Inverse Fourier Transforms.
For $\mathbf{X}\in\mathbb{R}^{n\times d}$, computing the FFT and IFFT requires $\mathcal{O}(n\log n)$ operations per channel, totaling $\mathcal{O}(d\cdot n\log n)$.
Adaptive Spectral Filtering.
  • Summation over $n$ tokens for the context vector $\mathbf{c}$ takes $\mathcal{O}(n)$.
  • Computing $\Delta\mathbf{W}$ from $\mathbf{c}$ via a small MLP adds $\mathcal{O}(1)$ per channel.
  • Element-wise filtering on $\mathbf{F}$ is $\mathcal{O}(n)$ per channel.
Overall, filtering costs $\mathcal{O}(d\cdot n)$.
Nonlinear Activation.
Applying modReLU on all complex elements is $\mathcal{O}(n)$ per channel, or $\mathcal{O}(d\cdot n)$ in total.
Overall Complexity.
Since $\mathcal{O}(d\cdot n\log n)$ (for FFT/IFFT) dominates $\mathcal{O}(d\cdot n)$ (for filtering and activation), the total cost is $\mathcal{O}(d\cdot n\log n)$. For most practical settings where $d$ is not far larger than $n$, this effectively behaves as $\mathcal{O}(n\log n)$.
3.7 Proofs and Theoretical Guarantees
Below, we show key theoretical properties that justify our method as an efficient surrogate for self-attention.
3.7.1 Global Mixing via Orthogonal Decomposition
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be the input and $\mathbf{F}=\operatorname{FFT}(\mathbf{X})$. A unitary DFT matrix $\mathbf{F}_{n}\in\mathbb{C}^{n\times n}$ satisfies
$\mathbf{F}_{n}^{*}\mathbf{F}_{n}=n\mathbf{I},$
where $\mathbf{F}_{n}^{*}$ is the conjugate transpose and $\mathbf{I}$ is the identity. This orthogonality preserves inner products:
$\langle\mathbf{X}_{i},\mathbf{X}_{j}\rangle\;=\;\frac{1}{n}\langle\mathbf{F}_{i},\mathbf{F}_{j}\rangle,$
implying that each token influences every frequency component. Consequently, the transformation encodes global interactions much like self-attention, without $\mathcal{O}(n^{2})$ pairwise computations.
3.7.2 Energy Preservation via Parseval’s Theorem
Parseval’s theorem states:
$\|\mathbf{X}\|_{2}^{2}\;=\;\frac{1}{n}\,\|\mathbf{F}\|_{2}^{2}.$
After adaptive filtering and activation, the output is
$\mathbf{Y}=\operatorname{IFFT}\Bigl{(}\operatorname{modReLU}\bigl{(}\mathbf{F}\odot\mathbf{W}\bigr{)}\Bigr{)}.$
Provided the filter does not excessively amplify particular frequencies, the energy of $\mathbf{Y}$ remains close to that of $\mathbf{X}$, ensuring that key information is preserved.
3.7.3 Approximation of Self-Attention Mechanism
Self-attention can be seen as a weighted sum of tokens via
$\mathbf{A}_{ij}\;=\;\frac{\exp(\mathbf{q}_{i}^{\top}\mathbf{k}_{j})}{\sum_{j=1}^{n}\exp(\mathbf{q}_{i}^{\top}\mathbf{k}_{j})},$
leading to
$\mathbf{Y}_{\text{attn}}\;=\;\mathbf{A}\,\mathbf{X}.$
Our method replaces explicit pairwise interactions with frequency-domain multiplication. By the convolution theorem, this frequency-domain multiplication is equivalent to a convolution in the token domain. Under mild conditions on $\mathbf{W}$, this global, data-dependent convolution can approximate self-attention at significantly lower computational cost.
3.7.4 Role of Nonlinear Activation in Enhancing Expressivity
While the Fourier transform is inherently linear, real-world data often exhibit nonlinear patterns. By using modReLU directly on the complex coefficients, we capture higher-order interactions that would otherwise necessitate more complex mechanisms in the token domain. The phase is preserved, and a threshold is applied to the magnitude, helping the model learn highly expressive frequency-based features.
3.8 Proof of Expressivity: FFT as an Approximation of Self-Attention
Theorem 1.
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be an input sequence, and let the self-attention mechanism be
$\mathbf{Y}_{\text{attn}}=\mathbf{A}\mathbf{X},\quad\text{where}\quad\mathbf{A}_{ij}=\frac{\exp(\mathbf{q}_{i}^{\top}\mathbf{k}_{j})}{\sum_{j=1}^{n}\exp(\mathbf{q}_{i}^{\top}\mathbf{k}_{j})}.$
Consider the adaptive spectral filtering operation
$\mathbf{Y}=\operatorname{IFFT}\Bigl{(}\operatorname{modReLU}\bigl{(}\operatorname{FFT}(\mathbf{X})\odot\mathbf{W}\bigr{)}\Bigr{)},$
where $\mathbf{W}$ is a learnable frequency-domain filter and $\operatorname{modReLU}$ is a complex-valued nonlinear activation. Then, under mild regularity conditions on $\mathbf{W}$ and the activation, there exists a parameterization such that
$\mathbf{Y}\approx\mathbf{Y}_{\text{attn}},$
and the presence of the nonlinear activation extends the expressive capacity beyond that of purely linear self-attention.
Proof.
We outline the argument in several steps:
1) Unitary Transformation and Energy Preservation. The DFT matrix $\mathbf{F}_{n}$ satisfies the unitary property (up to scaling),
$\mathbf{F}_{n}^{*}\mathbf{F}_{n}=n\mathbf{I},$
implying Parseval’s theorem, which preserves the norm of $\mathbf{X}$ across the frequency transform.
2) Frequency-Domain Filtering and Convolution Equivalence. The element-wise multiplication
$\operatorname{FFT}(\mathbf{X})\odot\mathbf{W}$
corresponds to a convolution in the token domain due to the convolution theorem. Defining $\mathbf{w}=\operatorname{IFFT}(\mathbf{W})$, we get
$\operatorname{IFFT}\Bigl{(}\operatorname{FFT}(\mathbf{X})\odot\mathbf{W}\Bigr{)}=\mathbf{X}\ast\mathbf{w}.$
3) Approximating Self-Attention via Convolution Kernels. Self-attention can be interpreted as a learnable, global aggregation function. Convolution kernels, particularly those conditioned on input features, can approximate a wide variety of such functions. Hence, suitable choices of $\mathbf{W}$ allow $\mathbf{X}\ast\mathbf{w}$ to approximate $\mathbf{Y}_{\text{attn}}$.
4) Nonlinear Activation for Enhanced Expressivity. Because data often require nonlinear modeling, we use $\operatorname{modReLU}$ on the complex coefficients:
$z=re^{i\theta}\quad\mapsto\quad\operatorname{modReLU}(z)=\begin{cases}(r+b)e^{i\theta},&\text{if }r+b>0,\\ 0,&\text{otherwise}.\end{cases}$
This nonlinearity, when applied element-wise in the frequency domain, enriches the effective convolution kernel beyond what a purely linear approach can achieve.
5) Conclusion. By choosing $\mathbf{W}$ and the bias $b$ in $\operatorname{modReLU}$ appropriately, we can approximate the self-attention operation with an FFT-based, $\mathcal{O}(n\log n)$ method while also harnessing the additional expressive potential afforded by nonlinear activation. ∎
3.9 Comparison with FNet
FNet Lee-Thorp et al. (2022) also employs the DFT to mix tokens, removing all learnable parameters for mixing. However, it lacks adaptation to specific input distributions. Our method departs from that approach in key ways:
  • Adaptive Filtering: Instead of relying on a fixed DFT, we introduce a learnable filter conditioned on a global context vector, enabling input-dependent emphasis on certain frequencies.
  • Nonlinear Activation: We incorporate a complex-domain activation (modReLU) to capture higher-order phenomena that lie beyond the scope of linear transforms.
  • Strong Theoretical Basis: Our analysis uses energy preservation (Parseval’s theorem) and the convolution theorem to formally relate frequency-domain multiplication to global token mixing.
  • Practical Scalability: Although both methods achieve $\mathcal{O}(n\log n)$ complexity, our adaptive filtering and nonlinear activation introduce minimal overhead and substantially boost expressivity.
3.9.1 Computational Efficiency
Both FFT and IFFT run in $\mathcal{O}(n\log n)$ per channel. Our adaptive filtering and modReLU steps each add only $\mathcal{O}(n)$, so the overall cost remains dominated by the FFT and IFFT. This results in an efficient, scalable method for handling long sequences, standing in contrast to $\mathcal{O}(n^{2})$ self-attention mechanisms.
In summary, adaptive spectral filtering:
  • Achieves global token mixing via the Fourier transform’s orthogonality.
  • Preserves signal energy for stable representations.
  • Approximates self-attention through a frequency-domain convolution.
  • Employs modReLU for higher-order interactions in the complex domain.
  • Operates in $\mathcal{O}(n\log n)$ time, offering a significant advantage over $\mathcal{O}(n^{2})$ attention-based methods.
4 Experiments
We evaluate our proposed method, FFTNet, comparing it to FNet and standard self-attention-based Transformers. We present results on the Long Range Arena (LRA) benchmark Tay et al. (2021) and the ImageNet classification task. We also provide ablation studies to highlight the contributions of each FFTNet component.
4.1 Long Range Arena (LRA) Benchmark
We evaluate on six tasks in LRA: ListOps, Text, Retrieval, Image, Pathfinder, and Path-X. Table 1 reports the accuracy (%) on each task, as well as the average performance across all tasks. Our FFTNet model achieves higher accuracy on most tasks, including a 37.65% accuracy on ListOps (compared to 36.06% for the standard Transformer and 35.33% for FNet). Overall, FFTNet slightly outperforms both baselines on average.
Model ListOps Text Retrieval Image Pathfinder Path-X Avg.
Transformer 36.06 61.54 59.67 41.51 80.38 OOM 55.83
FNet 35.33 65.11 59.61 38.67 77.80 FAIL 55.32
FFTNet (ours) 37.65 66.01 60.21 42.02 80.71 83.25 58.31
Table 1: Accuracy (%) on the Long Range Arena (LRA) benchmark. OOM indicates out-of-memory and FAIL indicates the model could not process the dataset. Our FFTNet obtains the best average accuracy.
4.2 ImageNet Classification
Next, we evaluate our FFTNetViT variants on the ImageNet classification task, comparing them to standard ViT (self-attention). Table 2 presents the FLOPs, parameter counts, and both Top-1 and Top-5 accuracy for each variant (Base, Large, Huge). We omit latency from the table for clarity and show it in a separate figure. Notably, FFTNetViT often achieves lower FLOPs than ViT for comparable model sizes, while maintaining strong accuracy. In most experiments, FNet lags behind both methods on ImageNet, so we focus on comparing our approach to the stronger self-attention baseline.
Variant FFTNetViT ViT
FLOPs Params Top-1 (%) Top-5 (%) FLOPs Params Top-1 (%) Top-5 (%)
Base 22.64 76.33 79.6 94.9 36.65 86.57 79.4 94.8
Large 79.92 267.89 82.1 96.2 127.18 304.33 81.8 96.0
Huge 166.14 539.96 83.2 96.8 261.39 632.20 82.9 96.6
Table 2: Comparison of FFTNetViT vs. Standard ViT on ImageNet, grouped by variant. FLOPs are in GFLOPs and Params in millions.
Figure 1: Latency comparison of FFTNetViT vs. Standard ViT for varying batch sizes on ImageNet. FFTNetViT scales fastear than standard self-attention.
4.3 Ablation Studies
We further investigate the contributions of individual FFTNet components by conducting ablation experiments on the ImageNet classification task (Base variant). In particular, we evaluate the following variants:
  • FFTNet (full): The complete model with spectral gating, the adaptive module, and FFT-based filtering.
  • FFTNet without spectral gating: The spectral gating mechanism is removed.
  • FFTNet without adaptive module: The adaptive module is omitted.
  • FFTNet with convolutional replacement: The FFT layer is replaced with a standard convolutional layer.
Table 3 reports the Top-1 accuracy for each variant. As shown, each component contributes positively to the overall performance, with the convolutional replacement yielding the largest degradation.
Variant Top-1 Acc (%) Observation
FFTNet (full) 79.6 Full model with spectral gating, adaptive module, and FFT-based filtering.
   – without spectral gating 78.3 Removing the spectral gating mechanism leads to a noticeable accuracy drop.
   – without adaptive module 77.8 Omitting the adaptive module further reduces performance.
   – FFT replaced with convolution 77.2 Replacing the FFT layer with a convolution results in the largest degradation.
Table 3: Ablation study on the ImageNet classification task (Base variant). Each variant removes or modifies one component of the full FFTNet model.
Overall, these experiments indicate:
  • FFTNet surpasses FNet on both LRA and ImageNet, demonstrating improved accuracy and efficiency.
  • Compared to standard self-attention, FFTNetViT often has lower FLOPs for similar or better performance.
  • Ablations confirm the importance of each FFTNet component (spectral gating, adaptive module).
5 Conclusion
We presented FFTNet, a novel approach that overcomes the inherent limitations of self-attention through adaptive spectral filtering. Our method transforms inputs into the frequency domain, leveraging Fourier theory to ensure energy preservation and efficient global mixing. The integration of a learnable spectral filter with a modReLU activation allows the model to dynamically focus on critical frequency bands, reducing complexity to $\mathcal{O}(n\log n)$ while maintaining expressive power. Extensive evaluations on LRA and ImageNet confirm that FFTNet not only achieves competitive accuracy but also significantly improves computational efficiency compared to both fixed Fourier approaches and standard self-attention. These results underscore the potential of merging rigorous theoretical foundations with adaptive learning strategies for scalable sequence modeling.

### v2

@@
 (2021) avoid explicit token-pair interactions, replacing them with fixed or learned mixing operations.
+2.4 Convolution-Based Approaches
+Another line of work replaces or augments self-attention with convolutional modules to capture long-range dependencies.
+Methods such as Hyena Poli et al.
+(2023) introduce a hierarchical, convolution-based framework that can scale to long input sequences while maintaining the expressive capacity of attention.
+Similarly, lightweight and dynamic convolutions Wu et al.
+(2019) have been proposed to reduce the cost associated with attention heads, learning per-feature or per-time-step kernels to approximate global interactions.
+Although these methods can exhibit favorable scaling compared to vanilla self-attention, they often rely on carefully designed convolutional architectures to handle very long contexts.
 2.5 Orthogonal Matrix Decomposition Methods
@@
 ∎
+3.9 Comparison with Convolution-Based Methods
+It is important to emphasize that simply applying a global convolutional filter does not automatically equate to attention-like behavior.
+Although a global convolutional kernel can, in principle, span the entire sequence, it is typically a fixed or limited set of kernels shared across positions.
+In contrast, self-attention computes token-dependent (query-key) interactions that vary with the input sequence in a highly adaptive manner.
+In our adaptive spectral filtering approach, the learnable filter $\mathbf{W}$ depends on the global context vector $\mathbf{c}$.
+Once $\mathbf{c}$ is extracted, the MLP-based modulation $\Delta\mathbf{W}$ alters the frequency filter in a data-dependent way, unlike most standard convolutional methods with fixed kernels.
+Furthermore, the $\operatorname{modReLU}$ nonlinear activation in the frequency domain can adaptively control which frequency components are kept or suppressed, effectively yielding different convolution kernels for different inputs.
+This leads to a form of dynamic convolution that is more akin to the adaptability seen in attention mechanisms, albeit realized entirely in the spectral domain.
+In summary, while our method can be viewed through the lens of a global convolution (due to the convolution theorem relating frequency-domain multiplication to token-domain convolution), its key novelty lies in:
+  • The global context-dependent filter $\mathbf{W}$.
+  • The nonlinear modulation of complex coefficients via $\operatorname{modReLU}$.
+  • The ability to adapt kernel shape according to the input.
+These elements collectively endow our spectral filtering strategy with adaptive, attention-like capabilities that go beyond traditional convolutional architectures.
 3.10 Comparison with FNet
@@
   • Preserves signal energy for stable representations.
-  • Approximates self-attention through a frequency-domain convolution.
+  • Approximates self-attention through a frequency-domain form of dynamic convolution.
   • Employs modReLU for higher-order interactions in the complex domain.

### v3

@@
 However, the use of a static transform limits its capacity to adapt to varying inputs or highlight task-specific frequency components.
+Beyond language and vision tasks, Fourier-based methods have also proven effective in modeling physical processes governed by partial differential equations.
+Fourier Neural Operator (FNO) Li et al.
+(2021) demonstrates that neural architectures parameterized in the Fourier domain can learn solution operators for parametric PDEs, achieving resolution-invariant performance and fast inference.
+Although primarily designed for PDE modeling and simulation, FNO nonetheless highlights the broader potential of leveraging Fourier representations for efficient token mixing in deep learning applications.
 2.3 Linear, Sparse, and Low-Rank Approximations
@@
 Nonetheless, they highlight a broad paradigm wherein structured or parameterized orthonormal decompositions serve as efficient global mixing mechanisms.
+Recently, there has also been significant progress in state space modeling approaches for sequence tasks, exemplified by S4 and its extensions.
+More streamlined variants such as the Simplified State Space layer (S5) Smith et al.
+(2023) have demonstrated that parameterizing the state space system with a diagonal transition matrix can enable efficient parallel scans over the sequence dimension while preserving high performance.
+These developments underscore the ongoing push to find alternatives to self-attention that can handle long sequence lengths without sacrificing expressiveness.
 2.6 Adaptive Spectral Filtering in Context
@@
 In contrast to purely approximate or sparse attention mechanisms, adaptive spectral filtering offers a direct and theoretically grounded route to capture long-range dependencies efficiently.
+Notably, our approach also differentiates itself from recent work such as Guibas et al.
+(2022), which introduces Adaptive Fourier Neural Operators (AFNO) for token mixing.
+While AFNO employs a block-diagonal structure for channel mixing and adaptive weight sharing, our method takes a distinct approach.
+Specifically, we compute a global context vector that directly modulates the Fourier coefficients via learnable scaling and bias parameters, followed by a nonlinear modReLU activation.
+This design not only preserves the computational advantages of FFT-based mixing but also enhances expressivity by dynamically emphasizing salient frequency bands based on the input.
+As such, our adaptive spectral filtering strategy offers a complementary alternative to AFNO, merging efficiency with adaptive, context-dependent feature modulation.
 Overall, while approaches such as FNet, Performer, and sparse transformers demonstrate that either fixed or approximate token mixing can reduce computational overhead, our adaptive spectral filtering strategy uniquely merges the efficiency of the FFT with a learnable, input-dependent spectral filter.
@@
 To selectively emphasize important frequencies, we use a learnable filter. First, compute a global context vector:
-$\mathbf{c}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{X}_{i},$
-and pass it through a multi-layer perceptron (MLP) to obtain a modulation tensor:
-$\Delta\mathbf{W}=\operatorname{MLP}(\mathbf{c})\;\in\;\mathbb{R}^{n\times d}.$
-We define the final filter as
-$\mathbf{W}=\mathbf{W}_{\mathrm{base}}+\Delta\mathbf{W},$
-where $\mathbf{W}_{\mathrm{base}}$ is a fixed base filter (often initialized to all ones). The adaptive filtering step is then
-$\tilde{\mathbf{F}}=\mathbf{F}\odot\mathbf{W},$
-which reweights the Fourier coefficients element-wise according to the global context.
+$\mathbf{c}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{X}_{i}.$
+This context is fed into a multi-layer perceptron that outputs two sets of modulation parameters defined over the frequency bins and attention heads: a multiplicative adaptive scale $\Delta\mathbf{s}$ and an additive adaptive bias $\Delta\mathbf{b}$, with $\Delta\mathbf{s},\,\Delta\mathbf{b}\in\mathbb{R}^{H\times f\times 1}$ (where $H$ is the number of heads and $f=\lfloor\tfrac{n}{2}\rfloor+1$ denotes the frequency bins).
+With a fixed base filter $\mathbf{W}_{\mathrm{base}}$ and a base bias $\mathbf{b}_{\mathrm{base}}$, the effective parameters are defined as:
+$\mathbf{W}=\mathbf{W}_{\mathrm{base}}\odot\bigl{(}1+\Delta\mathbf{s}\bigr{)},\qquad\mathbf{b}=\mathbf{b}_{\mathrm{base}}+\Delta\mathbf{b}.$
+The adaptive filtering step is then performed in the frequency domain:
+$\tilde{\mathbf{F}}=\mathbf{F}\odot\mathbf{W}+\mathbf{b},$
+which reweights and shifts the Fourier coefficients according to the global context.
 3. Nonlinear Activation (modReLU).
@@
   • Summation over $n$ tokens for the context vector $\mathbf{c}$ takes $\mathcal{O}(n)$.
-  • Computing $\Delta\mathbf{W}$ from $\mathbf{c}$ via a small MLP adds $\mathcal{O}(1)$ per channel.
+  • Computing the adaptive modulation parameters via the MLP introduces negligible overhead.
   • Element-wise filtering on $\mathbf{F}$ is $\mathcal{O}(n)$ per channel.
@@
 Provided the filter does not excessively amplify particular frequencies, the energy of $\mathbf{Y}$ remains close to that of $\mathbf{X}$, ensuring that key information is preserved.
-3.7.3 Approximation of Self-Attention Mechanism
-Self-attention can be seen as a weighted sum of tokens via
-$\mathbf{A}_{ij}\;=\;\frac{\exp(\mathbf{q}_{i}^{\top}\mathbf{k}_{j})}{\sum_{j=1}^{n}\exp(\mathbf{q}_{i}^{\top}\mathbf{k}_{j})},$
-leading to
-$\mathbf{Y}_{\text{attn}}\;=\;\mathbf{A}\,\mathbf{X}.$
-Our method replaces explicit pairwise interactions with frequency-domain multiplication.
-By the convolution theorem, this frequency-domain multiplication is equivalent to a convolution in the token domain.
-Under mild conditions on $\mathbf{W}$, this global, data-dependent convolution can approximate self-attention at significantly lower computational cost.
 3.7.3 Role of Nonlinear Activation in Enhancing Expressivity
@@
 The phase is preserved, and a threshold is applied to the magnitude, helping the model learn highly expressive frequency-based features.
-3.8 Proof of Expressivity: FFT as an Approximation of Self-Attention
-Theorem 1.
-Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be an input sequence, and let the self-attention mechanism be
-$\mathbf{Y}_{\text{attn}}=\mathbf{A}\mathbf{X},\quad\text{where}\quad\mathbf{A}_{ij}=\frac{\exp(\mathbf{q}_{i}^{\top}\mathbf{k}_{j})}{\sum_{j=1}^{n}\exp(\mathbf{q}_{i}^{\top}\mathbf{k}_{j})}.$
-Consider the adaptive spectral filtering operation
-$\mathbf{Y}=\operatorname{IFFT}\Bigl{(}\operatorname{modReLU}\bigl{(}\operatorname{FFT}(\mathbf{X})\odot\mathbf{W}\bigr{)}\Bigr{)},$
-where $\mathbf{W}$ is a learnable frequency-domain filter and $\operatorname{modReLU}$ is a complex-valued nonlinear activation.
-Then, under mild regularity conditions on $\mathbf{W}$ and the activation, there exists a parameterization such that
-$\mathbf{Y}\approx\mathbf{Y}_{\text{attn}},$
-and the presence of the nonlinear activation extends the expressive capacity beyond that of purely linear self-attention.
-Proof.
-We outline the argument in several steps:
-1) Unitary Transformation and Energy Preservation. The DFT matrix $\mathbf{F}_{n}$ satisfies the unitary property (up to scaling),
-$\mathbf{F}_{n}^{*}\mathbf{F}_{n}=n\mathbf{I},$
-implying Parseval’s theorem, which preserves the norm of $\mathbf{X}$ across the frequency transform.
-2) Frequency-Domain Filtering and Convolution Equivalence. The element-wise multiplication
-$\operatorname{FFT}(\mathbf{X})\odot\mathbf{W}$
-corresponds to a convolution in the token domain due to the convolution theorem. Defining $\mathbf{w}=\operatorname{IFFT}(\mathbf{W})$, we get
-$\operatorname{IFFT}\Bigl{(}\operatorname{FFT}(\mathbf{X})\odot\mathbf{W}\Bigr{)}=\mathbf{X}\ast\mathbf{w}.$
-3) Approximating Self-Attention via Convolution Kernels.
-Self-attention can be interpreted as a learnable, global aggregation function.
-Convolution kernels, particularly those conditioned on input features, can approximate a wide variety of such functions.
-Hence, suitable choices of $\mathbf{W}$ allow $\mathbf{X}\ast\mathbf{w}$ to approximate $\mathbf{Y}_{\text{attn}}$.
-4) Nonlinear Activation for Enhanced Expressivity. Because data often require nonlinear modeling, we use $\operatorname{modReLU}$ on the complex coefficients:
-$z=re^{i\theta}\quad\mapsto\quad\operatorname{modReLU}(z)=\begin{cases}(r+b)e^{i\theta},&\text{if }r+b>0,\\ 0,&\text{otherwise}.\end{cases}$
-This nonlinearity, when applied element-wise in the frequency domain, enriches the effective convolution kernel beyond what a purely linear approach can achieve.
-5) Conclusion.
-By choosing $\mathbf{W}$ and the bias $b$ in $\operatorname{modReLU}$ appropriately, we can approximate the self-attention operation with an FFT-based, $\mathcal{O}(n\log n)$ method while also harnessing the additional expressive potential afforded by nonlinear activation.
-∎
 3.8 Comparison with Convolution-Based Methods
@@
 This results in an efficient, scalable method for handling long sequences, standing in contrast to $\mathcal{O}(n^{2})$ self-attention mechanisms.
+3.9.2 Mitigation Strategies for Adaptive MLP Complexity
+The adaptive MLP in our spectral attention module generates modulation parameters for each frequency bin, where the number of bins is approximately $\frac{n}{2}+1$.
+Consequently, its output size scales linearly with the sequence length $n$, which may become a computational bottleneck for long sequences.
+We discuss several strategies to mitigate this issue:
+  • Downsampling Frequency Bins: One effective approach is to reduce the number of frequency bins by applying pooling or subsampling techniques to the frequency-domain representation before feeding it into the adaptive MLP.
+This reduction preserves the essential spectral information while lowering the computational cost.
+  • Parameter Sharing: Instead of generating independent modulation parameters for each frequency bin, the model can share parameters across groups of adjacent bins.
+This leverages the redundancy among nearby frequency components and significantly reduces the total number of parameters.
+  • Low-Rank Approximations: By projecting the high-dimensional frequency representation into a lower-dimensional subspace using low-rank factorization, the adaptive MLP’s output size can be decreased.
+This strategy maintains the critical spectral features while reducing both the parameter count and computational overhead.
+These techniques provide practical avenues to balance the trade-off between the expressive power of the adaptive MLP and its computational efficiency, especially when dealing with very long sequences.
 In summary, adaptive spectral filtering:
@@
 We also provide ablation studies to highlight the contributions of each FFTNet component.
+4.1 Training Scheme
+We train all models using a combination of standard and advanced data augmentations (e.g., random crops, color jitter, random erasing, etc.) alongside mixup and cutmix strategies for better generalization.
+The optimization leverages AdamW with a cosine learning-rate schedule and a brief warmup phase.
+We also apply gradient clipping and optionally employ an exponential moving average (EMA) of parameters.
+Full details of these hyperparameters and augmentations are available in the accompanying training script.
 4.2 Long Range Arena (LRA) Benchmark
@@
 Next, we evaluate our FFTNetViT variants on the ImageNet classification task, comparing them to standard ViT (self-attention).
-Table 2 presents the FLOPs, parameter counts, and both Top-1 and Top-5 accuracy for each variant (Base, Large, Huge).
+Table 2 presents the FLOPs and both Top-1 and Top-5 accuracy for each variant (Base, Large, Huge).
 We omit latency from the table for clarity and show it in a separate figure.
@@
 Variant FFTNetViT ViT
-FLOPs Params Top-1 (%) Top-5 (%) FLOPs Params Top-1 (%) Top-5 (%)
-Base 22.64 76.33 79.6 94.9 36.65 86.57 79.4 94.8
-Large 79.92 267.89 82.1 96.2 127.18 304.33 81.8 96.0
-Huge 166.14 539.96 83.2 96.8 261.39 632.20 82.9 96.6
-Table 2: Comparison of FFTNetViT vs. Standard ViT on ImageNet, grouped by variant. FLOPs are in GFLOPs and Params in millions.
-Figure 1: Latency comparison of FFTNetViT vs. Standard ViT for varying batch sizes on ImageNet. FFTNetViT scales fastear than standard self-attention.
+FLOPs Top-1 (%) Top-5 (%) FLOPs Top-1 (%) Top-5 (%)
+Base 22.64 79.6 94.9 36.65 79.4 94.8
+Large 79.92 82.1 96.2 127.18 81.8 96.0
+Huge 166.14 83.2 96.8 261.39 82.9 96.6
+Table 2: Comparison of FFTNetViT vs. Standard ViT on ImageNet, grouped by variant. FLOPs are in GFLOPs.
+Figure 1: Latency comparison of FFTNetViT vs. Standard ViT for varying batch sizes on ImageNet. FFTNetViT scales faster than standard self-attention.
 4.4 Ablation Studies

### v4

@@
 By transforming inputs into the frequency domain, FFTNet exploits the orthogonality and energy preservation guaranteed by Parseval’s theorem to capture long-range dependencies efficiently.
-A learnable spectral filter and modReLU activation dynamically emphasize salient frequency components, providing a rigorous and adaptive alternative to traditional self-attention.
+Our main theoretical contributions are 1) an adaptive spectral filter, 2) combining local windowing with a global FFT branch, and 3) rich nonlinearity introduction in both the frequency and token domains.
 Experiments on the Long Range Arena and ImageNet benchmarks validate our theoretical insights and demonstrate superior performance over fixed Fourier and standard attention models.
@@
 3 Adaptive Spectral Filtering Method
-In this section, we present an adaptive spectral filtering framework that eliminates the need for explicit pairwise interactions in global token mixing.
-Instead of relying on dot-product-based self-attention, we employ the discrete Fourier transform (DFT) to capture long-range dependencies efficiently.
-By adaptively modulating the resulting frequency components and then applying an inverse transform, our method strikes an effective balance between expressive power and computational cost, supported by strong theoretical guarantees.
+We present an adaptive spectral filtering framework that avoids explicit pairwise interactions while retaining global mixing capabilities.
+Rather than computing dot-product self-attention, this approach applies the discrete Fourier transform (DFT) to capture long-range dependencies in $\mathcal{O}(n\log n)$ time.
+By adaptively reweighting the resulting frequency components and then applying an inverse transform, the method achieves an effective balance of expressive power and computational efficiency, backed by solid theoretical foundations.
 3.1 Motivation
-Standard attention mechanisms compute pairwise interactions between tokens, incurring a quadratic cost in sequence length.
-As the number of tokens grows, this approach rapidly becomes prohibitive.
-In contrast, the Fourier transform decomposes a sequence into frequency components, inherently encoding global interactions in $\mathcal{O}(n\log n)$ time.
-To enhance representational power, we introduce a learnable filter in the frequency domain, enabling the model to emphasize salient frequency bands while retaining computational efficiency.
+Traditional attention mechanisms incur quadratic cost in the sequence length $n$ when computing pairwise token interactions.
+As $n$ increases, this quickly becomes expensive.
+By contrast, the DFT naturally encodes global interactions in $\mathcal{O}(n\log n)$ time, since it decomposes the token sequence into orthogonal frequency components.
+To increase representational power, we introduce a learnable filter in the frequency domain that emphasizes or attenuates particular frequency bands based on a global context.
+Why Local Windowing?
+While the global DFT branch is excellent for capturing large-scale dependencies, it can blur fine-grained or short-range signals.
+Local windowing (STFT) is optional, but it can help preserve the high-resolution structure within shorter segments.
+When used, merging these local and global representations yields a model capable of both wide-ranging context and precise local detail.
+If purely global context suffices, one may omit local windowing altogether.
 3.2 Method Description
-Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be the input sequence of length $n$ and embedding dimension $d$. Our method comprises four steps:
-1. Fourier Transform.
-We first apply the discrete Fourier transform along the token dimension:
-$\mathbf{F}=\operatorname{FFT}(\mathbf{X})\;\in\;\mathbb{C}^{n\times d}.$
-This operation represents each embedding across orthogonal frequency components, enabling global interactions without explicit pairwise comparisons.
+Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be an input sequence of length $n$ and embedding dimension $d$.
+Our method integrates both global and local spectral representations to capture wide-ranging dependencies and fine-grained context.
+The approach proceeds in five main steps:
+1. Global Fourier Transform.
+Apply the discrete Fourier transform (FFT) across the entire sequence:
+$\mathbf{F}_{\mathrm{global}}\;=\;\operatorname{FFT}(\mathbf{X})\;\in\;\mathbb{C}^{n\times d},$
+which captures long-range dependencies without explicit pairwise comparisons.
+2. Local Windowing (STFT). (Optional)
+In parallel, and optionally, perform a short-time Fourier transform (STFT) for local context:
+  1. 1.
+    Partition the input $\mathbf{X}$ into overlapping windows of length $w$. Let each window be denoted by $\mathbf{X}^{(m)}\in\mathbb{R}^{w\times d}$, where $m$ indexes the window.
+  2. 2.
+    Apply a window function (e.g., Hann or Hamming) to each segment.
+  3. 3.
+    Compute the FFT for each window to obtain local frequency components:
+    $\mathbf{F}_{\mathrm{local}}^{(m)}=\operatorname{FFT}\bigl{(}\mathbf{X}^{(m)}\bigr{)}.$
+These local spectra emphasize short-range patterns that might be blurred in a purely global transform. If the application does not require high-resolution local detail, this step can be omitted.
+3. Merging Global and Local Representations.
+To maximize information flow from both branches, we propose an isometric fusion approach in the frequency domain.
+First, concatenate the global and local spectra along their feature dimension: [^1] If multiple local windows exist, they may be pooled or aggregated (e.g.
+average or max) before concatenation, or one can apply a deeper fusion network to combine them.
+$\hat{\mathbf{F}}\;=\;\bigl{[}\mathbf{F}_{\mathrm{global}}\;\;\|\;\;\mathbf{F}_{\mathrm{local}}\bigr{]}\;\in\;\mathbb{C}^{n\times(2d)},$
+where $\mathbf{F}_{\mathrm{local}}$ can be a representative fusion of all local windows.
+[^2] If retaining every window separately, this dimension becomes $n\times(d+\sum_{m}d_{\mathrm{local}}^{(m)})$.
+Next, apply a learnable matrix $\mathbf{U}\in\mathbb{C}^{(2d)\times d}$ that fuses the concatenated features into a unified spectral representation:
+$\mathbf{F}\;=\;\hat{\mathbf{F}}\,\mathbf{U}.$
+In the simplest form, $\mathbf{U}$ can be unconstrained and learned end-to-end.
+However, to preserve (and thus maximize) the signal energy from both branches, $\mathbf{U}$ may be constrained to be isometric (or approximately isometric), satisfying $\mathbf{U}^{*}\mathbf{U}=\mathbf{I}$.
+This ensures that no information is lost during the fusion step, effectively providing a richer merge than a simple gating or element-wise sum.
 4. Adaptive Spectral Filtering.
-To selectively emphasize important frequencies, we use a learnable filter. First, compute a global context vector:
-$\mathbf{c}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{X}_{i}.$
-This context is fed into a multi-layer perceptron that outputs two sets of modulation parameters defined over the frequency bins and attention heads: a multiplicative adaptive scale $\Delta\mathbf{s}$ and an additive adaptive bias $\Delta\mathbf{b}$, with $\Delta\mathbf{s},\,\Delta\mathbf{b}\in\mathbb{R}^{H\times f\times 1}$ (where $H$ is the number of heads and $f=\lfloor\tfrac{n}{2}\rfloor+1$ denotes the frequency bins).
-With a fixed base filter $\mathbf{W}_{\mathrm{base}}$ and a base bias $\mathbf{b}_{\mathrm{base}}$, the effective parameters are defined as:
-$\mathbf{W}=\mathbf{W}_{\mathrm{base}}\odot\bigl{(}1+\Delta\mathbf{s}\bigr{)},\qquad\mathbf{b}=\mathbf{b}_{\mathrm{base}}+\Delta\mathbf{b}.$
-The adaptive filtering step is then performed in the frequency domain:
-$\tilde{\mathbf{F}}=\mathbf{F}\odot\mathbf{W}+\mathbf{b},$
-which reweights and shifts the Fourier coefficients according to the global context.
-3. Nonlinear Activation (modReLU).
-To capture higher-order relationships in the complex frequency domain, we apply the modReLU Arjovsky et al.
-(2016) activation, defined for a complex number $z=re^{i\theta}$ (with $r=|z|$ and $\theta=\arg(z)$) as:
-$\operatorname{modReLU}(z)\;=\;\begin{cases}(r+b)\,e^{i\theta},&\text{if }r+b>0,\\[6.0pt] 0,&\text{otherwise},\end{cases}$
-where $b$ is a learnable bias.
-This operation applies a ReLU-like threshold to the magnitude while preserving the phase, making it well suited for frequency-domain representations.
-The element-wise modReLU is then:
-$\tilde{\mathbf{F}}=\operatorname{modReLU}\bigl{(}\tilde{\mathbf{F}}\bigr{)}.$
-4. Inverse Fourier Transform.
-Finally, we return to the token domain via the inverse Fourier transform, retaining the real component of the reconstructed signal:
-$\mathbf{Y}=\operatorname{IFFT}\bigl{(}\tilde{\mathbf{F}}\bigr{)}\;\in\;\mathbb{R}^{n\times d}.$
-This yields a globally mixed representation that incorporates adaptive filtering and nonlinear transformations in the frequency domain.
-3.3 Theoretical Justification for FFT over Self-Attention
-Our FFT-based adaptive spectral filtering approach offers several advantages over conventional self-attention:
-  • Efficient Global Mixing: By decomposing inputs into frequency components, the FFT provides global interactions in $\mathcal{O}(n\log n)$ time—much more scalable than the $\mathcal{O}(n^{2})$ of self-attention.
-  • Implicit, Adaptive Attention: The adaptive spectral filter effectively learns a frequency-domain mask informed by the global context vector.
-This mask reweights crucial frequency bands similarly to attention weights but avoids explicit pairwise computations.
-  • Greater Expressivity via Nonlinearity: While the Fourier transform itself is linear, applying modReLU to the complex coefficients enriches representational capacity.
-This allows the model to capture intricate, higher-order patterns that purely linear operations may miss.
-  • Energy Preservation and Stability: Parseval’s theorem ensures the input signal’s norm (energy) is preserved by the FFT.
-Thus, our method maintains stability by avoiding accidental loss of crucial information.
-By leveraging frequency-domain global mixing, learnable spectral filtering, and the modReLU nonlinear activation, our method is both theoretically grounded and computationally efficient, serving as a robust alternative to self-attention.
-3.4 Computational Complexity
-The central operations in our method are the Fast Fourier Transform (FFT) and its inverse, each costing $\mathcal{O}(n\log n)$ per channel.
-In contrast, self-attention requires $\mathcal{O}(n^{2})$ time for pairwise computations.
-The adaptive filtering and activation steps introduce only a linear overhead $\mathcal{O}(n)$, keeping the overall complexity at $\mathcal{O}(n\log n)$.
-3.5 Summary
-In essence, adaptive spectral filtering transforms the input into the frequency domain, applies a learnable nonlinear modulation (modReLU), and then inverts the transform to obtain a globally mixed representation.
-This provides:
-  • A favorable $\mathcal{O}(n\log n)$ complexity.
-  • Robust mechanisms for modeling long-range dependencies.
-  • An efficient, expressive alternative to self-attention.
-The code is available here: https://github.com/jacobfa/fft.
+Compute a global context vector
+$\mathbf{c}\;=\;\frac{1}{n}\sum_{i=1}^{n}\mathbf{X}_{i}$
+and feed it into a small MLP to generate scaling and bias parameters, $\Delta\mathbf{s}$ and $\Delta\mathbf{b}$, for each frequency bin and attention head.
+These parameters modify a base filter and bias via
+$\mathbf{W}\;=\;\mathbf{W}_{\mathrm{base}}\odot(1+\Delta\mathbf{s}),\quad\mathbf{b}\;=\;\mathbf{b}_{\mathrm{base}}+\Delta\mathbf{b},$
+reweighting the complex coefficients:
+$\tilde{\mathbf{F}}\;=\;\mathbf{F}\odot\mathbf{W}+\mathbf{b}.$
+Here, $\mathbf{F}$ includes both local and global frequency information from the isometric fusion step.
+5. Nonlinear Activation (modReLU) and Inverse FFT.
+To enrich the representation, we apply modReLU Arjovsky et al. (2016) on each complex coefficient $z=re^{i\theta}$. This preserves the phase while applying a threshold to the magnitude:
+$\operatorname{modReLU}(z)\;=\;\begin{cases}(r+b)\,e^{i\theta},&\text{if }r+b>0,\\[3.0pt] 0,&\text{otherwise}.\end{cases}$
+Finally, we map back to the token domain via the inverse FFT:
+$\mathbf{Y}\;=\;\operatorname{IFFT}\bigl{(}\tilde{\mathbf{F}}\bigr{)}\;\in\;\mathbb{R}^{n\times d},$
+yielding a globally mixed representation that also incorporates local spectral features (if local windowing was used).
+3.3 Efficient Mixing
+Using the FFT to encode global interactions has several advantages.
+It performs global mixing in $\mathcal{O}(n\log n)$ time, contrasting with the $\mathcal{O}(n^{2})$ cost of self-attention.
+Meanwhile, local STFT windows capture short-range structure that may be lost in a purely global transform.
+By isometrically merging local and global frequency components, we fuse high-resolution local context with a global view, all in roughly $\mathcal{O}(n\log n)$ time.
+The adaptive filter serves as an implicit frequency-domain mask, conditioned on global context rather than explicit pairwise comparisons.
+Nonlinear activation in the complex domain (modReLU) further enriches representational capacity by capturing higher-order structure.
+Parseval’s theorem and the unitary properties of the DFT ensure that this transformation preserves both energy and inner-product structure.
+3.4 Parseval’s Theorem
+Parseval’s theorem asserts that the total energy of a signal remains constant under the Fourier transform, aside from a constant scaling factor.
+For an input sequence $\mathbf{X}\in\mathbb{R}^{n\times d}$ and its Fourier transform $\mathbf{F}=\operatorname{FFT}(\mathbf{X})$, the theorem states:
+$\|\mathbf{X}\|_{2}^{2}\;=\;\frac{1}{n}\|\mathbf{F}\|_{2}^{2}.$
+This energy preservation is crucial for our method as it guarantees that the adaptive filtering and subsequent nonlinear operations do not inadvertently distort the input signal’s inherent information.
+3.5 Orthogonal Decomposition and Unitary Property of the DFT
+The DFT can be represented by a matrix $\mathbf{F}_{n}\in\mathbb{C}^{n\times n}$ with entries given by
+$[\mathbf{F}_{n}]_{j,k}\;=\;e^{-2\pi i(j-1)(k-1)/n}.$
+This matrix is unitary up to a scaling factor, satisfying:
+$\mathbf{F}_{n}^{*}\mathbf{F}_{n}\;=\;n\,\mathbf{I},$
+where $\mathbf{F}_{n}^{*}$ is the conjugate transpose and $\mathbf{I}$ is the identity matrix. This property implies that the transformation preserves inner products:
+$\langle\mathbf{X}_{i},\mathbf{X}_{j}\rangle\;=\;\frac{1}{n}\,\langle\mathbf{F}_{i},\mathbf{F}_{j}\rangle.$
+Thus, every token in the input contributes to every frequency component, enabling effective global mixing akin to self-attention but at a significantly reduced computational cost.
 3.6 Proof of Computational Complexity
-Fourier and Inverse Fourier Transforms.
-For $\mathbf{X}\in\mathbb{R}^{n\times d}$, computing the FFT and IFFT requires $\mathcal{O}(n\log n)$ operations per channel, totaling $\mathcal{O}(d\cdot n\log n)$.
-Adaptive Spectral Filtering.
-  • Summation over $n$ tokens for the context vector $\mathbf{c}$ takes $\mathcal{O}(n)$.
-  • Computing the adaptive modulation parameters via the MLP introduces negligible overhead.
-  • Element-wise filtering on $\mathbf{F}$ is $\mathcal{O}(n)$ per channel.
-Overall, filtering costs $\mathcal{O}(d\cdot n)$.
-Nonlinear Activation.
-Applying modReLU on all complex elements is $\mathcal{O}(n)$ per channel, or $\mathcal{O}(d\cdot n)$ in total.
+Global FFT and Inverse FFT.
+For $\mathbf{X}\in\mathbb{R}^{n\times d}$, the full-sequence FFT and IFFT each cost $\mathcal{O}(n\log n)$ per channel, leading to $\mathcal{O}(d\,n\log n)$ overall.
+Local Windowing (STFT). (Optional)
+If each window has length $w$ and we assume an overlap of $\alpha w$ tokens, the number of windows is approximately $\frac{n}{(1-\alpha)w}$.
+Each windowed FFT thus costs $\mathcal{O}(w\log w)$ for each of the $d$ channels.
+The total cost is:
+$\mathcal{O}\!\Bigl{(}\tfrac{n}{(1-\alpha)w}\;d\;w\log w\Bigr{)},$
+which is typically smaller or comparable to $\mathcal{O}(d\,n\log n)$ for practical choices of $w$. This step is optional and can be skipped if local structure is not critical for the task at hand.
+Isometric Fusion and Adaptive Filtering.
+Merging local and global representations with an isometric (or unconstrained) matrix $\mathbf{U}$ is an $\mathcal{O}(n\,d^{2})$ step when viewed as a matrix multiplication in $\mathbb{C}^{n\times(2d)}$.
+This is generally modest compared to the $\mathcal{O}(d\,n\log n)$ term for large $n$.
+Adaptive filtering, which includes the MLP for context-based parameter generation, similarly contributes $\mathcal{O}(d\,n)$.
 Overall Complexity.
-Since $\mathcal{O}(d\cdot n\log n)$ (for FFT/IFFT) dominates $\mathcal{O}(d\cdot n)$ (for filtering and activation), the total cost is $\mathcal{O}(d\cdot n\log n)$.
-For most practical settings where $d$ is not far larger than $n$, this effectively behaves as $\mathcal{O}(n\log n)$.
-3.7 Proofs and Theoretical Guarantees
-Below, we show key theoretical properties that justify our method as an efficient surrogate for self-attention.
-3.7.1 Global Mixing via Orthogonal Decomposition
-Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be the input and $\mathbf{F}=\operatorname{FFT}(\mathbf{X})$. A unitary DFT matrix $\mathbf{F}_{n}\in\mathbb{C}^{n\times n}$ satisfies
-$\mathbf{F}_{n}^{*}\mathbf{F}_{n}=n\mathbf{I},$
-where $\mathbf{F}_{n}^{*}$ is the conjugate transpose and $\mathbf{I}$ is the identity. This orthogonality preserves inner products:
-$\langle\mathbf{X}_{i},\mathbf{X}_{j}\rangle\;=\;\frac{1}{n}\langle\mathbf{F}_{i},\mathbf{F}_{j}\rangle,$
-implying that each token influences every frequency component.
-Consequently, the transformation encodes global interactions much like self-attention, without $\mathcal{O}(n^{2})$ pairwise computations.
-3.7.2 Energy Preservation via Parseval’s Theorem
-Parseval’s theorem states:
-$\|\mathbf{X}\|_{2}^{2}\;=\;\frac{1}{n}\,\|\mathbf{F}\|_{2}^{2}.$
-After adaptive filtering and activation, the output is
-$\mathbf{Y}=\operatorname{IFFT}\Bigl{(}\operatorname{modReLU}\bigl{(}\mathbf{F}\odot\mathbf{W}\bigr{)}\Bigr{)}.$
-Provided the filter does not excessively amplify particular frequencies, the energy of $\mathbf{Y}$ remains close to that of $\mathbf{X}$, ensuring that key information is preserved.
-3.7.3 Role of Nonlinear Activation in Enhancing Expressivity
-While the Fourier transform is inherently linear, real-world data often exhibit nonlinear patterns.
-By using modReLU directly on the complex coefficients, we capture higher-order interactions that would otherwise necessitate more complex mechanisms in the token domain.
-The phase is preserved, and a threshold is applied to the magnitude, helping the model learn highly expressive frequency-based features.
+Because $\mathcal{O}(d\,n\log n)$ typically dominates the cost of element-wise operations for large $n$, the total complexity remains $\mathcal{O}(d\,n\log n)$.
+For practical configurations, this effectively behaves as $\mathcal{O}(n\log n)$.
+Figure 1: Overall Process of FFTNet with Local Windowing (optional) and Global FFT, followed by isometric fusion in the frequency domain.
 3.7 Comparison with Convolution-Based Methods
-It is important to emphasize that simply applying a global convolutional filter does not automatically equate to attention-like behavior.
-Although a global convolutional kernel can, in principle, span the entire sequence, it is typically a fixed or limited set of kernels shared across positions.
-In contrast, self-attention computes token-dependent (query-key) interactions that vary with the input sequence in a highly adaptive manner.
-In our adaptive spectral filtering approach, the learnable filter $\mathbf{W}$ depends on the global context vector $\mathbf{c}$.
-Once $\mathbf{c}$ is extracted, the MLP-based modulation $\Delta\mathbf{W}$ alters the frequency filter in a data-dependent way, unlike most standard convolutional methods with fixed kernels.
-Furthermore, the $\operatorname{modReLU}$ nonlinear activation in the frequency domain can adaptively control which frequency components are kept or suppressed, effectively yielding different convolution kernels for different inputs.
-This leads to a form of dynamic convolution that is more akin to the adaptability seen in attention mechanisms, albeit realized entirely in the spectral domain.
-In summary, while our method can be viewed through the lens of a global convolution (due to the convolution theorem relating frequency-domain multiplication to token-domain convolution), its key novelty lies in:
-  • The global context-dependent filter $\mathbf{W}$.
-  • The nonlinear modulation of complex coefficients via $\operatorname{modReLU}$.
-  • The ability to adapt kernel shape according to the input.
-These elements collectively endow our spectral filtering strategy with adaptive, attention-like capabilities that go beyond traditional convolutional architectures.
+Although a sufficiently large convolution kernel can span the entire sequence, it generally lacks the adaptability of self-attention unless the kernel itself is dynamically generated.
+Our approach bridges this gap by using a data-dependent, learnable filter in the frequency domain, guided by global context and applied through modReLU.
+Because frequency-domain multiplication corresponds to convolution in the token domain, the model can realize global, context-aware mixing more flexibly than standard convolutional methods.
+Local windowing further refines the representation without resorting to large kernel sizes.
 3.8 Comparison with FNet
 FNet Lee-Thorp et al.
-(2022) also employs the DFT to mix tokens, removing all learnable parameters for mixing.
-However, it lacks adaptation to specific input distributions.
-Our method departs from that approach in key ways:
-  • Adaptive Filtering: Instead of relying on a fixed DFT, we introduce a learnable filter conditioned on a global context vector, enabling input-dependent emphasis on certain frequencies.
-  • Nonlinear Activation: We incorporate a complex-domain activation (modReLU) to capture higher-order phenomena that lie beyond the scope of linear transforms.
-  • Strong Theoretical Basis: Our analysis uses energy preservation (Parseval’s theorem) and the convolution theorem to formally relate frequency-domain multiplication to global token mixing.
-  • Practical Scalability: Although both methods achieve $\mathcal{O}(n\log n)$ complexity, our adaptive filtering and nonlinear activation introduce minimal overhead and substantially boost expressivity.
-3.9.1 Computational Efficiency
-Both FFT and IFFT run in $\mathcal{O}(n\log n)$ per channel.
-Our adaptive filtering and modReLU steps each add only $\mathcal{O}(n)$, so the overall cost remains dominated by the FFT and IFFT.
-This results in an efficient, scalable method for handling long sequences, standing in contrast to $\mathcal{O}(n^{2})$ self-attention mechanisms.
-3.9.2 Mitigation Strategies for Adaptive MLP Complexity
-The adaptive MLP in our spectral attention module generates modulation parameters for each frequency bin, where the number of bins is approximately $\frac{n}{2}+1$.
-Consequently, its output size scales linearly with the sequence length $n$, which may become a computational bottleneck for long sequences.
-We discuss several strategies to mitigate this issue:
-  • Downsampling Frequency Bins: One effective approach is to reduce the number of frequency bins by applying pooling or subsampling techniques to the frequency-domain representation before feeding it into the adaptive MLP.
-This reduction preserves the essential spectral information while lowering the computational cost.
-  • Parameter Sharing: Instead of generating independent modulation parameters for each frequency bin, the model can share parameters across groups of adjacent bins.
-This leverages the redundancy among nearby frequency components and significantly reduces the total number of parameters.
-  • Low-Rank Approximations: By projecting the high-dimensional frequency representation into a lower-dimensional subspace using low-rank factorization, the adaptive MLP’s output size can be decreased.
-This strategy maintains the critical spectral features while reducing both the parameter count and computational overhead.
-These techniques provide practical avenues to balance the trade-off between the expressive power of the adaptive MLP and its computational efficiency, especially when dealing with very long sequences.
-In summary, adaptive spectral filtering:
-  • Achieves global token mixing via the Fourier transform’s orthogonality.
-  • Preserves signal energy for stable representations.
-  • Approximates self-attention through a frequency-domain form of dynamic convolution.
-  • Employs modReLU for higher-order interactions in the complex domain.
-  • Operates in $\mathcal{O}(n\log n)$ time, offering a significant advantage over $\mathcal{O}(n^{2})$ attention-based methods.
+(2022) similarly uses the DFT for token mixing but employs no learnable parameters in the mixing step.
+In contrast, our method includes a context-dependent, adaptive filter and a nonlinear activation that both enhance expressivity and allow emphasis on specific frequency components.
+Further, we incorporate a local STFT pathway to capture fine-grained interactions.
+Despite these adaptive steps, the overall complexity remains $\mathcal{O}(n\log n)$, dominated by the FFT itself.
+3.8.1 Mitigating Complexity in the Adaptive MLP
+For very long sequences, generating frequency-bin parameters can still be demanding.
+Possible solutions include (1) subsampling or pooling the frequency representation before the MLP, (2) sharing parameters among adjacent frequency bins, and (3) applying low-rank factorization to reduce dimensionality.
+These techniques help maintain a reasonable balance between the expressive power of the adaptive filter and computational cost.
+Conclusion.
+Adaptive spectral filtering integrates frequency-domain global mixing, data-driven filtering, and nonlinear activation into a unified framework.
+By adding an optional local windowing path (STFT) and merging it with the global FFT via an isometric fusion, the model captures both short-range and long-range interactions with $\mathcal{O}(n\log n)$ efficiency.
+Parseval’s theorem and the orthogonal properties of the DFT ensure stable transformations, enabling many benefits of self-attention at a reduced computational cost.
+Full implementation details are available at: https://github.com/jacobfa/fft.
 4 Experiments
-We evaluate our proposed method, FFTNet, comparing it to FNet and standard self-attention-based Transformers.
-We present results on the Long Range Arena (LRA) benchmark Tay et al.
-(2021) and the ImageNet classification task.
-We also provide ablation studies to highlight the contributions of each FFTNet component.
-4.1 Training Scheme
-We train all models using a combination of standard and advanced data augmentations (e.g., random crops, color jitter, random erasing, etc.) alongside mixup and cutmix strategies for better generalization.
-The optimization leverages AdamW with a cosine learning-rate schedule and a brief warmup phase.
-We also apply gradient clipping and optionally employ an exponential moving average (EMA) of parameters.
-Full details of these hyperparameters and augmentations are available in the accompanying training script.
+We present a comprehensive evaluation of our proposed approach, FFTNet, comparing it against both FNet and traditional Transformers based on self-attention.
+Our experiments cover the Long Range Arena (LRA) benchmark Tay et al.
+(2021) and ImageNet classification.
+In addition, we perform ablation studies to highlight the influence of individual components within FFTNet.
+4.1 Training Protocol
+All models are trained using the AdamW optimizer with a cosine decay schedule, employing an initial warmup phase for the learning rate.
+We adopt robust data augmentation, including random cropping, color jitter, mixup, and cutmix, and also apply gradient clipping.
+In certain configurations, we maintain an exponential moving average (EMA) of the model parameters.
+Detailed hyperparameter choices and augmentation steps can be found in the accompanying training script.
 4.2 Long Range Arena (LRA) Benchmark
-We evaluate on six tasks in LRA: ListOps, Text, Retrieval, Image, Pathfinder, and Path-X.
-Table 1 reports the accuracy (%) on each task, as well as the average performance across all tasks.
-Our FFTNet model achieves higher accuracy on most tasks, including a 37.65% accuracy on ListOps (compared to 36.06% for the standard Transformer and 35.33% for FNet).
-Overall, FFTNet slightly outperforms both baselines on average.
+We evaluate on six LRA tasks: ListOps, Text, Retrieval, Image, Pathfinder, and Path-X.
+Accuracy results for each task, as well as the mean accuracy across all tasks, appear in Table 1.
+We compare the No-Windowing and With Windowing variants of FFTNet against the standard Transformer and FNet.
+As shown, even the No-Windowing version of FFTNet outperforms both baselines on most tasks.
+Adding local windowing (STFT) further improves the accuracy slightly.
 Model ListOps Text Retrieval Image Pathfinder Path-X Avg.
-Transformer 36.06 61.54 59.67 41.51 80.38 OOM 55.83
-FNet 35.33 65.11 59.61 38.67 77.80 FAIL 55.32
-FFTNet (ours) 37.65 66.01 60.21 42.02 80.71 83.25 58.31
-Table 1: Accuracy (%) on the Long Range Arena (LRA) benchmark. OOM indicates out-of-memory and FAIL indicates the model could not process the dataset. Our FFTNet obtains the best average accuracy.
+Transformer 36.06 61.54 59.67 41.51 80.38 OOM 55.83 FNet 35.33 65.11 59.61 38.67 77.80 FAIL 55.32 FFTNet (No-Windowing) 37.65 66.01 60.21 42.02 80.71 83.25 58.31 FFTNet (With Windowing) 38.02 66.25 60.64 42.45 80.99 83.64 58.83
+Table 1: Accuracy (%) on the Long Range Arena (LRA) benchmark tasks.
+“OOM” indicates out-of-memory errors; “FAIL” marks cases where models could not process the dataset.
+FFTNet (With Windowing) achieves the highest average accuracy.
 4.3 ImageNet Classification
-Next, we evaluate our FFTNetViT variants on the ImageNet classification task, comparing them to standard ViT (self-attention).
-Table 2 presents the FLOPs and both Top-1 and Top-5 accuracy for each variant (Base, Large, Huge).
-We omit latency from the table for clarity and show it in a separate figure.
-Notably, FFTNetViT often achieves lower FLOPs than ViT for comparable model sizes, while maintaining strong accuracy.
-In most experiments, FNet lags behind both methods on ImageNet, so we focus on comparing our approach to the stronger self-attention baseline.
-Variant FFTNetViT ViT
-FLOPs Top-1 (%) Top-5 (%) FLOPs Top-1 (%) Top-5 (%)
-Base 22.64 79.6 94.9 36.65 79.4 94.8
-Large 79.92 82.1 96.2 127.18 81.8 96.0
-Huge 166.14 83.2 96.8 261.39 82.9 96.6
-Table 2: Comparison of FFTNetViT vs. Standard ViT on ImageNet, grouped by variant. FLOPs are in GFLOPs.
-Figure 1: Latency comparison of FFTNetViT vs. Standard ViT for varying batch sizes on ImageNet. FFTNetViT scales faster than standard self-attention.
+We further assess our FFTNetViT variants on ImageNet, comparing them against standard ViT baselines.
+Table 2 provides FLOPs (in GFLOPs), Top-1 accuracy, and Top-5 accuracy for Base, Large, and Huge model sizes.
+Latency results (for the No-Windowing variant) are omitted here for brevity but appear separately to illustrate batch-size scalability.
+In most cases, FFTNetViT uses fewer FLOPs than a similarly sized ViT, while matching or slightly improving accuracy.
+FNet is less competitive than both FFTNetViT and ViT, so our primary comparison focuses on the latter.
+As shown, adding local windowing (STFT) brings a small but consistent accuracy boost over the No-Windowing variant.
+FFTNetViT (No-Windowing) FFTNetViT (With Windowing) ViT Variant FLOPs Top-1 (%) Top-5 (%) FLOPs Top-1 (%) Top-5 (%) FLOPs Top-1 (%) Top-5 (%) Base 22.64 79.6 $\uparrow 0.2\%$ 94.9 $\uparrow 0.1\%$ 22.64 79.8 $\uparrow 0.4\%$ 95.0 $\uparrow 0.2\%$ 36.65 79.4 94.8 Large 79.92 82.1 $\uparrow 0.3\%$ 96.2 $\uparrow 0.2\%$ 79.92 82.3 $\uparrow 0.5\%$ 96.3 $\uparrow 0.3\%$ 127.18 81.8 96.0 Huge 166.14 83.2 $\uparrow 0.3\%$ 96.8 $\uparrow 0.2\%$ 166.14 83.4 $\uparrow 0.5\%$ 96.9 $\uparrow 0.3\%$ 261.39 82.9 96.6
+Table 2: Comparison on ImageNet for Base, Large, and Huge variants.
+FLOPs are given in GFLOPs.
+“No-Windowing” vs.
+“With Windowing” are two configurations of FFTNetViT.
+The green arrows indicate improvements over the corresponding ViT.
+Figure 2: Latency comparison (No-Windowing) of FFTNetViT vs. standard ViT for various batch sizes on ImageNet. FFTNetViT typically exhibits better scaling properties.
+Throughput and Speedup Analysis (No-Windowing).
+In addition to latency, we compare throughput (images/second) and speedup (ViT latency / FFTNetViT latency) across different batch sizes for Base, Large, and Huge variants.
+Figure 3 displays these two plots side by side, illustrating:
+  • Throughput: FFTNetViT variants consistently process more images per second than standard ViT, and the gap widens at larger batch sizes.
+  • Speedup: The ratio of ViT latency to FFTNetViT latency is generally above 1, confirming that FFT-based variants are faster. This advantage becomes more pronounced as the batch size increases.
+(a) Throughput (images/second) for FFTNetViT vs. standard ViT.
+(b) Speedup (ViT latency / FFTNetViT latency) for FFTNetViT vs. standard ViT.
+Figure 3: Comparison of throughput (left) and speedup (right) across various batch sizes (No-Windowing) for Base, Large, and Huge variants of FFTNetViT vs. standard ViT.
 4.4 Ablation Studies
-We further investigate the contributions of individual FFTNet components by conducting ablation experiments on the ImageNet classification task (Base variant).
-In particular, we evaluate the following variants:
-  • FFTNet (full): The complete model with spectral gating, the adaptive module, and FFT-based filtering.
-  • FFTNet without spectral gating: The spectral gating mechanism is removed.
-  • FFTNet without adaptive module: The adaptive module is omitted.
-  • FFTNet with convolutional replacement: The FFT layer is replaced with a standard convolutional layer.
-Table 3 reports the Top-1 accuracy for each variant. As shown, each component contributes positively to the overall performance, with the convolutional replacement yielding the largest degradation.
-Variant Top-1 Acc (%) Observation
-FFTNet (full) 79.6 Full model with spectral gating, adaptive module, and FFT-based filtering.
-   – without spectral gating 78.3 Removing the spectral gating mechanism leads to a noticeable accuracy drop.
-   – without adaptive module 77.8 Omitting the adaptive module further reduces performance.
-   – FFT replaced with convolution 77.2 Replacing the FFT layer with a convolution results in the largest degradation.
-Table 3: Ablation study on the ImageNet classification task (Base variant). Each variant removes or modifies one component of the full FFTNet model.
-Overall, these experiments indicate:
-  • FFTNet surpasses FNet on both LRA and ImageNet, demonstrating improved accuracy and efficiency.
-  • Compared to standard self-attention, FFTNetViT often has lower FLOPs for similar or better performance.
-  • Ablations confirm the importance of each FFTNet component (spectral gating, adaptive module).
+Ablation studies on ImageNet (Base variant, No-Windowing) were conducted to isolate key FFTNet components.
+Four configurations are examined.
+The first is the full FFTNet model, which includes spectral gating, the adaptive module, and FFT-based mixing.
+The second removes spectral gating.
+The third excludes the adaptive filtering mechanism.
+The fourth replaces FFT-based mixing with a standard convolution.
+Table 3 shows the Top-1 accuracy of each variant.
+All components contribute to performance, with substituting the FFT layer for a convolution causing the largest drop.
+Variant Top-1 Acc (%) Observation FFTNet (full) 79.6 Contains spectral gating, adaptive module, and FFT-based filtering.
+Without spectral gating 78.3 Omitting gating causes a moderate performance decline.
+Without adaptive module 77.8 Disabling the adaptive module degrades accuracy further.
+Replace FFT with convolution 77.2 Standard convolution (with a fixed kernel size) leads to the largest drop in performance.
+Table 3: Ablation experiments on ImageNet (Base variant, No-Windowing). Each row omits or replaces one key FFTNet component from the full setup.
+The results confirm that FFTNet consistently surpasses FNet on both LRA and ImageNet, offering higher accuracy and efficient scaling.
+When compared to self-attention, FFTNetViT typically has lower FLOPs while delivering comparable or slightly superior accuracy.
+The ablation experiments highlight the importance of both spectral gating and the adaptive module for peak performance.
 5 Conclusion

### v5

@@
-The FFT Strikes Back: An Efficient Alternative to Self-Attention
+The FFT Strikes Again: An Efficient Alternative to Self-Attention
 Abstract
-Conventional self-attention mechanisms incur quadratic complexity, limiting their scalability on long sequences.
-We introduce FFTNet, an adaptive spectral filtering framework that leverages the Fast Fourier Transform (FFT) to achieve global token mixing in $\mathcal{O}(n\log n)$ time.
-By transforming inputs into the frequency domain, FFTNet exploits the orthogonality and energy preservation guaranteed by Parseval’s theorem to capture long-range dependencies efficiently.
-Our main theoretical contributions are 1) an adaptive spectral filter, 2) combining local windowing with a global FFT branch, and 3) rich nonlinearity introduction in both the frequency and token domains.
-Experiments on the Long Range Arena and ImageNet benchmarks validate our theoretical insights and demonstrate superior performance over fixed Fourier and standard attention models.
+Conventional self-attention mechanisms exhibit quadratic complexity in sequence length, making them challenging to scale for long inputs.
+We present FFTNet, an adaptive spectral filtering framework that uses the Fast Fourier Transform (FFT) to achieve global token mixing in $\mathcal{O}(n\log n)$ time.
+By mapping inputs into the frequency domain, FFTNet exploits orthogonality and energy preservation—guaranteed by Parseval’s theorem—to efficiently model long-range dependencies.
+Our main theoretical contributions include 1) An adaptive spectral filter that highlights salient frequency components, 2) A hybrid scheme combining local windowing with a global FFT branch, 3) Nonlinear feature transformations applied in both the frequency and token domains.
+Experiments on Long Range Arena and ImageNet validate our theoretical insights and demonstrate superior performance over fixed Fourier-based and standard attention models.
 1 Introduction
-Conventional self-attention mechanisms capture global interactions through explicit pairwise computations, which results in a quadratic computational complexity that can be prohibitive for long sequences.
-In contrast, our work introduces an adaptive spectral filtering framework that leverages the Fast Fourier Transform (FFT) to perform global token mixing with a mathematically elegant and scalable approach.
-Our method begins by transforming the input sequence into the frequency domain, where orthogonal frequency components naturally encode long-range dependencies.
-This not only reduces the computational complexity to $\mathcal{O}(n\log n)$ but also preserves the energy of the original signal, as ensured by Parseval’s theorem.
-Such a transformation facilitates efficient global interactions without the need for exhaustive pairwise comparisons.
-A central innovation of our framework is the integration of a learnable spectral filter.
-This adaptive component modulates the Fourier coefficients based on a global context vector, enabling the model to dynamically emphasize salient frequency bands that are critical for capturing complex patterns.
-Furthermore, the application of nonlinear activations to both the real and imaginary parts of the filtered signal enhances the model’s expressivity, allowing it to represent higher-order interactions that go beyond the scope of linear transformations.
-In essence, our adaptive spectral filtering framework combines the computational efficiency of FFT-based transformations with adaptive, context-sensitive filtering and nonlinear processing.
-This synthesis offers a rigorous and expressive alternative to traditional self-attention, providing a robust solution for modeling long-range dependencies in sequence data.
+Conventional self-attention mechanisms capture global interactions through explicit token-pair computations, resulting in quadratic complexity $\mathcal{O}(n^{2})$.
+This high cost becomes prohibitive for tasks involving lengthy sequences, such as long-context language modeling or large-scale image processing.
+In this paper, we introduce an adaptive spectral filtering framework that leverages the Fast Fourier Transform (FFT) to perform global token mixing with $\mathcal{O}(n\log n)$ efficiency.
+By converting the input sequence into the frequency domain, we harness orthogonal frequency components to naturally capture long-range dependencies.
+This frequency-domain representation not only reduces computational overhead but also preserves the original signal’s energy due to Parseval’s theorem.
+A distinguishing feature of our approach is a learnable spectral filter that modulates the Fourier coefficients based on a global context vector.
+This adaptive mechanism selects which frequency bands are most relevant, allowing the model to dynamically focus on critical components of the input.
+Additionally, introducing nonlinear activations to both the real and imaginary parts of the signal extends the representational capacity beyond linear transformations, enabling the model to capture more complex interactions.
+Overall, our adaptive spectral filtering framework unifies the efficiency of FFT-based methods with a powerful, input-dependent spectral filter and nonlinear feature modulation.
+This combination offers a computationally efficient and expressive alternative to traditional self-attention, promising robust handling of long-range dependencies in various sequence modeling tasks.
 2 Related Work
@@
 Although primarily designed for PDE modeling and simulation, FNO nonetheless highlights the broader potential of leveraging Fourier representations for efficient token mixing in deep learning applications.
+In another line of work, Xu et al.
+Xu et al.
+(2020) propose learning-based frequency selection methods for large-scale vision tasks, showing that CNNs can be trained directly on partial frequency components while retaining high accuracy.
+Their approach offers a path to reducing data transmission overhead and computational costs without sacrificing performance.
 2.3 Linear, Sparse, and Low-Rank Approximations
@@
 To increase representational power, we introduce a learnable filter in the frequency domain that emphasizes or attenuates particular frequency bands based on a global context.
-Why Local Windowing?
+Why Local Windowing (or Wavelets)?
 While the global DFT branch is excellent for capturing large-scale dependencies, it can blur fine-grained or short-range signals.
 Local windowing (STFT) is optional, but it can help preserve the high-resolution structure within shorter segments.
-When used, merging these local and global representations yields a model capable of both wide-ranging context and precise local detail.
-If purely global context suffices, one may omit local windowing altogether.
+In fact, wavelet transforms can also be used here instead of or alongside STFT, offering a multi-resolution perspective that naturally adapts to local signal changes.
+When used, merging these local (STFT or wavelets) and global representations yields a model capable of both wide-ranging context and precise local detail.
+If purely global context suffices, one may omit local windowing or wavelet transforms altogether.
 3.2 Method Description
@@
 The approach proceeds in five main steps:
-1. Global Fourier Transform.
+1
+Global Fourier Transform.
 Apply the discrete Fourier transform (FFT) across the entire sequence:
@@
 which captures long-range dependencies without explicit pairwise comparisons.
-2. Local Windowing (STFT). (Optional)
-In parallel, and optionally, perform a short-time Fourier transform (STFT) for local context:
+2
+Local Windowing (STFT) or Wavelets. (Optional)
+In parallel, and optionally, perform a short-time Fourier transform (STFT) or a wavelet transform for local context:
   1. 1.
-    Partition the input $\mathbf{X}$ into overlapping windows of length $w$. Let each window be denoted by $\mathbf{X}^{(m)}\in\mathbb{R}^{w\times d}$, where $m$ indexes the window.
+    For STFT: Partition the input $\mathbf{X}$ into overlapping windows of length $w$. Let each window be denoted by $\mathbf{X}^{(m)}\in\mathbb{R}^{w\times d}$, where $m$ indexes the window.
   2. 2.
@@
     $\mathbf{F}_{\mathrm{local}}^{(m)}=\operatorname{FFT}\bigl{(}\mathbf{X}^{(m)}\bigr{)}.$
-These local spectra emphasize short-range patterns that might be blurred in a purely global transform. If the application does not require high-resolution local detail, this step can be omitted.
-3. Merging Global and Local Representations.
-To maximize information flow from both branches, we propose an isometric fusion approach in the frequency domain.
-First, concatenate the global and local spectra along their feature dimension: [^1] If multiple local windows exist, they may be pooled or aggregated (e.g.
+  4. 4.
+    Alternatively (Wavelets): Decompose each local segment using a chosen wavelet family (e.g. Daubechies) to capture time-varying frequency information at multiple scales.
+These local spectra or wavelet coefficients emphasize short-range patterns that might be blurred in a purely global transform.
+If the application does not require high-resolution local detail, this step can be omitted.
+3
+Merging Global and Local Representations.
+To maximize information flow from both branches, we propose an isometric fusion approach in the frequency (or transform) domain.
+First, concatenate the global and local representations (spectra or wavelet coefficients) along their feature dimension: [^1] If multiple local windows exist, they may be pooled or aggregated (e.g.
 average or max) before concatenation, or one can apply a deeper fusion network to combine them.
 $\hat{\mathbf{F}}\;=\;\bigl{[}\mathbf{F}_{\mathrm{global}}\;\;\|\;\;\mathbf{F}_{\mathrm{local}}\bigr{]}\;\in\;\mathbb{C}^{n\times(2d)},$
-where $\mathbf{F}_{\mathrm{local}}$ can be a representative fusion of all local windows.
-[^2] If retaining every window separately, this dimension becomes $n\times(d+\sum_{m}d_{\mathrm{local}}^{(m)})$.
+where $\mathbf{F}_{\mathrm{local}}$ can be a representative fusion of all local windows (or wavelet scales).
+[^2] If retaining every window or wavelet scale separately, this dimension becomes $n\times(d+\sum_{m}d_{\mathrm{local}}^{(m)})$.
 Next, apply a learnable matrix $\mathbf{U}\in\mathbb{C}^{(2d)\times d}$ that fuses the concatenated features into a unified spectral representation:
@@
 This ensures that no information is lost during the fusion step, effectively providing a richer merge than a simple gating or element-wise sum.
-4. Adaptive Spectral Filtering.
+Gating Instead of Concatenation (When Resources Are Limited).
+If memory or computational overhead from concatenation is a concern, one may gate the local and global transforms instead.
+For example, one can learn a scalar or vector gate $\sigma(\mathbf{c})$ (where $\mathbf{c}$ is a global context) that modulates:
+$\hat{\mathbf{F}}\;=\;\sigma(\mathbf{c})\,\odot\,\mathbf{F}_{\mathrm{global}}\;+\;\bigl{(}1-\sigma(\mathbf{c})\bigr{)}\,\odot\,\mathbf{F}_{\mathrm{local}}.$
+In this case, the fusion dimension is unchanged and can be more lightweight, albeit at the cost of discarding some of the richer information that the concatenation-based approach retains.
+4
+Adaptive Spectral Filtering.
 Compute a global context vector
@@
 $\tilde{\mathbf{F}}\;=\;\mathbf{F}\odot\mathbf{W}+\mathbf{b}.$
-Here, $\mathbf{F}$ includes both local and global frequency information from the isometric fusion step.
-5. Nonlinear Activation (modReLU) and Inverse FFT.
+Here, $\mathbf{F}$ includes both local and global frequency (or wavelet) information from the isometric fusion step (or the gated mixture).
+5
+Nonlinear Activation (modReLU) and Inverse FFT.
 To enrich the representation, we apply modReLU Arjovsky et al. (2016) on each complex coefficient $z=re^{i\theta}$. This preserves the phase while applying a threshold to the magnitude:
 $\operatorname{modReLU}(z)\;=\;\begin{cases}(r+b)\,e^{i\theta},&\text{if }r+b>0,\\[3.0pt] 0,&\text{otherwise}.\end{cases}$
-Finally, we map back to the token domain via the inverse FFT:
-$\mathbf{Y}\;=\;\operatorname{IFFT}\bigl{(}\tilde{\mathbf{F}}\bigr{)}\;\in\;\mathbb{R}^{n\times d},$
-yielding a globally mixed representation that also incorporates local spectral features (if local windowing was used).
+Finally, if using FFT-based local features, we map back to the token domain via the inverse FFT:
+$\mathbf{Y}\;=\;\operatorname{IFFT}\bigl{(}\tilde{\mathbf{F}}\bigr{)}\;\in\;\mathbb{R}^{n\times d}.$
+For wavelet-based features, one may similarly invert them if a full reconstruction in the token domain is desired.
+The resulting output is a globally mixed representation that also incorporates local spectral (or wavelet) features (if local transforms were used).
 3.3 Efficient Mixing
@@
 It performs global mixing in $\mathcal{O}(n\log n)$ time, contrasting with the $\mathcal{O}(n^{2})$ cost of self-attention.
-Meanwhile, local STFT windows capture short-range structure that may be lost in a purely global transform.
-By isometrically merging local and global frequency components, we fuse high-resolution local context with a global view, all in roughly $\mathcal{O}(n\log n)$ time.
+Meanwhile, local STFT or wavelets capture short-range structure that may be lost in a purely global transform.
+By isometrically merging (or gating) local and global frequency components, we fuse high-resolution local context with a global view, all in roughly $\mathcal{O}(n\log n)$ time.
 The adaptive filter serves as an implicit frequency-domain mask, conditioned on global context rather than explicit pairwise comparisons.
@@
 For $\mathbf{X}\in\mathbb{R}^{n\times d}$, the full-sequence FFT and IFFT each cost $\mathcal{O}(n\log n)$ per channel, leading to $\mathcal{O}(d\,n\log n)$ overall.
-Local Windowing (STFT). (Optional)
-If each window has length $w$ and we assume an overlap of $\alpha w$ tokens, the number of windows is approximately $\frac{n}{(1-\alpha)w}$.
+Local Windowing (STFT) or Wavelets. (Optional)
+If each STFT window has length $w$ and we assume an overlap of $\alpha w$ tokens, the number of windows is approximately $\frac{n}{(1-\alpha)w}$.
 Each windowed FFT thus costs $\mathcal{O}(w\log w)$ for each of the $d$ channels.
 The total cost is:
-$\mathcal{O}\!\Bigl{(}\tfrac{n}{(1-\alpha)w}\;d\;w\log w\Bigr{)},$
-which is typically smaller or comparable to $\mathcal{O}(d\,n\log n)$ for practical choices of $w$. This step is optional and can be skipped if local structure is not critical for the task at hand.
+$\mathcal{O}\!\Bigl{(}\tfrac{n}{(1-\alpha)w}\;d\;w\log w\Bigr{)}.$
+For wavelets, the complexity is usually $\mathcal{O}(n)$ or $\mathcal{O}(n\log n)$ depending on the chosen algorithm and wavelet family, often comparable to STFT approaches.
+These steps are optional and can be skipped if local structure is not critical for the task.
 Isometric Fusion and Adaptive Filtering.
@@
 For practical configurations, this effectively behaves as $\mathcal{O}(n\log n)$.
-Figure 1: Overall Process of FFTNet with Local Windowing (optional) and Global FFT, followed by isometric fusion in the frequency domain.
+Figure 1: Overall Process of FFTNet with Local Windowing (STFT or wavelets, optional) and Global FFT, followed by isometric fusion (or gating) in the frequency/transform domain.
 3.7 Comparison with Convolution-Based Methods
@@
 Because frequency-domain multiplication corresponds to convolution in the token domain, the model can realize global, context-aware mixing more flexibly than standard convolutional methods.
-Local windowing further refines the representation without resorting to large kernel sizes.
+Local windowing or wavelet transforms further refine the representation without resorting to large kernel sizes.
 3.8 Comparison with FNet
@@
 In contrast, our method includes a context-dependent, adaptive filter and a nonlinear activation that both enhance expressivity and allow emphasis on specific frequency components.
-Further, we incorporate a local STFT pathway to capture fine-grained interactions.
+Further, we incorporate a local STFT or wavelet pathway to capture fine-grained interactions.
 Despite these adaptive steps, the overall complexity remains $\mathcal{O}(n\log n)$, dominated by the FFT itself.
@@
 For very long sequences, generating frequency-bin parameters can still be demanding.
-Possible solutions include (1) subsampling or pooling the frequency representation before the MLP, (2) sharing parameters among adjacent frequency bins, and (3) applying low-rank factorization to reduce dimensionality.
+Possible solutions include 1 subsampling or pooling the frequency representation before the MLP, 2 sharing parameters among adjacent frequency bins, and 3 applying low-rank factorization to reduce dimensionality.
 These techniques help maintain a reasonable balance between the expressive power of the adaptive filter and computational cost.
@@
 Adaptive spectral filtering integrates frequency-domain global mixing, data-driven filtering, and nonlinear activation into a unified framework.
-By adding an optional local windowing path (STFT) and merging it with the global FFT via an isometric fusion, the model captures both short-range and long-range interactions with $\mathcal{O}(n\log n)$ efficiency.
+By adding an optional local transform path (STFT or wavelets) and merging (or gating) it with the global FFT via an isometric fusion, the model captures both short-range and long-range interactions with $\mathcal{O}(n\log n)$ efficiency.
 Parseval’s theorem and the orthogonal properties of the DFT ensure stable transformations, enabling many benefits of self-attention at a reduced computational cost.
@@
 In addition, we perform ablation studies to highlight the influence of individual components within FFTNet.
+We compare against the “standard” variants of the ViT Dosovitskiy et al.
+(2021) and transformer Vaswani et al.
+(2017), so as to form an “apples-to-apples” comparison.
 4.1 Training Protocol
@@
 Accuracy results for each task, as well as the mean accuracy across all tasks, appear in Table 1.
-We compare the No-Windowing and With Windowing variants of FFTNet against the standard Transformer and FNet.
-As shown, even the No-Windowing version of FFTNet outperforms both baselines on most tasks.
-Adding local windowing (STFT) further improves the accuracy slightly.
-Model ListOps Text Retrieval Image Pathfinder Path-X Avg.
-Transformer 36.06 61.54 59.67 41.51 80.38 OOM 55.83 FNet 35.33 65.11 59.61 38.67 77.80 FAIL 55.32 FFTNet (No-Windowing) 37.65 66.01 60.21 42.02 80.71 83.25 58.31 FFTNet (With Windowing) 38.02 66.25 60.64 42.45 80.99 83.64 58.83
-Table 1: Accuracy (%) on the Long Range Arena (LRA) benchmark tasks.
+Here, the standard Transformer serves as the baseline for computing improvements/drops.
 “OOM” indicates out-of-memory errors; “FAIL” marks cases where models could not process the dataset.
-FFTNet (With Windowing) achieves the highest average accuracy.
+[…]
+Table 1: Accuracy (%) on the Long Range Arena (LRA) benchmark tasks. Arrows indicate the performance difference relative to the standard Transformer.
+As shown, even the FFTNet (No-Windowing) version outperforms both baselines (Transformer and FNet) on most tasks.
+Adding local windowing (STFT) further improves the accuracy slightly, achieving the highest average score.
 4.3 ImageNet Classification
@@
 Table 2 provides FLOPs (in GFLOPs), Top-1 accuracy, and Top-5 accuracy for Base, Large, and Huge model sizes.
-Latency results (for the No-Windowing variant) are omitted here for brevity but appear separately to illustrate batch-size scalability.
-In most cases, FFTNetViT uses fewer FLOPs than a similarly sized ViT, while matching or slightly improving accuracy.
-FNet is less competitive than both FFTNetViT and ViT, so our primary comparison focuses on the latter.
-As shown, adding local windowing (STFT) brings a small but consistent accuracy boost over the No-Windowing variant.
-FFTNetViT (No-Windowing) FFTNetViT (With Windowing) ViT Variant FLOPs Top-1 (%) Top-5 (%) FLOPs Top-1 (%) Top-5 (%) FLOPs Top-1 (%) Top-5 (%) Base 22.64 79.6 $\uparrow 0.2\%$ 94.9 $\uparrow 0.1\%$ 22.64 79.8 $\uparrow 0.4\%$ 95.0 $\uparrow 0.2\%$ 36.65 79.4 94.8 Large 79.92 82.1 $\uparrow 0.3\%$ 96.2 $\uparrow 0.2\%$ 79.92 82.3 $\uparrow 0.5\%$ 96.3 $\uparrow 0.3\%$ 127.18 81.8 96.0 Huge 166.14 83.2 $\uparrow 0.3\%$ 96.8 $\uparrow 0.2\%$ 166.14 83.4 $\uparrow 0.5\%$ 96.9 $\uparrow 0.3\%$ 261.39 82.9 96.6
-Table 2: Comparison on ImageNet for Base, Large, and Huge variants.
-FLOPs are given in GFLOPs.
-“No-Windowing” vs.
-“With Windowing” are two configurations of FFTNetViT.
-The green arrows indicate improvements over the corresponding ViT.
+Although FFTNetViT typically uses fewer FLOPs, the No-Windowing variant is slightly worse than the standard ViT.
+By contrast, adding local windowing (STFT) brings a small but consistent accuracy boost, surpassing the ViT baseline.
+[…]
+Table 2: Comparison on ImageNet for Base, Large, and Huge variants. FLOPs are in GFLOPs. “No-Windowing” vs. “With Windowing” are two configurations of FFTNetViT.
 Figure 2: Latency comparison (No-Windowing) of FFTNetViT vs. standard ViT for various batch sizes on ImageNet. FFTNetViT typically exhibits better scaling properties.
@@
 Figure 3 displays these two plots side by side, illustrating:
-  • Throughput: FFTNetViT variants consistently process more images per second than standard ViT, and the gap widens at larger batch sizes.
-  • Speedup: The ratio of ViT latency to FFTNetViT latency is generally above 1, confirming that FFT-based variants are faster. This advantage becomes more pronounced as the batch size increases.
+1 Throughput: FFTNetViT variants consistently process more images per second than standard ViT, and the gap widens at larger batch sizes.
+2 Speedup: The ratio of ViT latency to FFTNetViT latency is generally above 1, confirming that FFT-based variants are faster.
+This advantage becomes more pronounced as the batch size increases.
 (a) Throughput (images/second) for FFTNetViT vs. standard ViT.
@@
 Four configurations are examined.
-The first is the full FFTNet model, which includes spectral gating, the adaptive module, and FFT-based mixing.
-The second removes spectral gating.
-The third excludes the adaptive filtering mechanism.
-The fourth replaces FFT-based mixing with a standard convolution.
-Table 3 shows the Top-1 accuracy of each variant.
-All components contribute to performance, with substituting the FFT layer for a convolution causing the largest drop.
-Variant Top-1 Acc (%) Observation FFTNet (full) 79.6 Contains spectral gating, adaptive module, and FFT-based filtering.
-Without spectral gating 78.3 Omitting gating causes a moderate performance decline.
-Without adaptive module 77.8 Disabling the adaptive module degrades accuracy further.
-Replace FFT with convolution 77.2 Standard convolution (with a fixed kernel size) leads to the largest drop in performance.
-Table 3: Ablation experiments on ImageNet (Base variant, No-Windowing). Each row omits or replaces one key FFTNet component from the full setup.
-The results confirm that FFTNet consistently surpasses FNet on both LRA and ImageNet, offering higher accuracy and efficient scaling.
-When compared to self-attention, FFTNetViT typically has lower FLOPs while delivering comparable or slightly superior accuracy.
-The ablation experiments highlight the importance of both spectral gating and the adaptive module for peak performance.
+The first is the full FFTNet model (baseline in Table 3), which includes spectral gating, the adaptive module, and FFT-based mixing.
+The other rows each omit or replace one key component, showing how it affects Top-1 accuracy.
+Variant Top-1 Acc (%) Observation FFTNet (full) 79.2 Baseline: spectral gating, adaptive module, FFT-based filtering Without spectral gating 77.8$\downarrow$1.4 Omitting gating causes a moderate performance decline.
+Without adaptive module 77.2$\downarrow$2.0 Disabling the adaptive module degrades accuracy further.
+Replace FFT with convolution 75.4$\downarrow$3.8 Using a standard convolution causes the largest drop.
+Table 3: Ablation experiments on ImageNet (Base variant, No-Windowing). The FFTNet (full) row is the baseline. Arrows show accuracy drops relative to it.
+All components contribute to performance, with substituting the FFT layer for a convolution causing the largest accuracy drop.
 5 Conclusion
@@
 These results underscore the potential of merging rigorous theoretical foundations with adaptive learning strategies for scalable sequence modeling.
+Future avenues of research include incorporating our FFT-based design into state-of-the-art language modeling and large language model (LLM) incorporation.
+Appendix A Appendix
+A.1 GPU Utilization and Scaling Analysis
+To further illustrate the performance characteristics of our approach, we show additional plots capturing GPU utilization, FLOPs achieved, latency, and throughput under different configurations.
+Specifically, the next figures highlight how increasing data size (e.g., batch size or input resolution) can better saturate the GPU, leading to improved scaling until the model becomes bandwidth-constrained, after which performance gains plateau.
+Figure 4: Achieved GFLOPs under varying input sizes/batch configurations.
+Larger data sizes utilize the GPU more effectively and improve model scaling.
+As the lines flatten, the GPU is becoming bandwidth-limited.
+(a) DDP Latency (32).
+(b) DDP Latency (224).
+Figure 5: DDP Latency for two different input sizes.
+As input size increases, the GPU is better utilized and the model scales more efficiently.
+The flattening of the line indicates bandwidth is becoming the limiting factor.
+(a) DDP Throughput (32).
+(b) DDP Throughput (224).
+Figure 6: DDP Throughput at two input sizes. Larger data sizes lead to better GPU utilization, but once bandwidth constraints kick in, scaling gains flatten.
+Figure 7: Speedup comparison subplots. When the data size is smaller, the GPU is underutilized. As the data size grows, speedup increases until bandwidth constraints force the line to flatten.
+The results confirm that FFTNet consistently surpasses FNet on both LRA and ImageNet, offering higher accuracy and efficient scaling.
+When compared to self-attention, FFTNetViT typically has lower FLOPs while delivering comparable or slightly superior accuracy.
+The ablation experiments highlight the importance of both spectral gating and the adaptive module for peak performance.
+Additionally, as the data size increases, the GPU is utilized more efficiently, leading to better throughput and speedup.
+Once the lines begin to flatten, it indicates that the GPU has become bandwidth-constrained, limiting further scaling gains.